### v1

# Distributional Instruments: Identification and Estimation with Quantile Least Squares

Abstract
We study instrumental-variable designs where policy reforms strongly shift the distribution of an endogenous variable but only weakly move its mean. We formalize this by introducing distributional relevance: instruments may be “purely distributional" with $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ while $F_{X\mid Z}(\cdot\mid Z)$ varies nontrivially. Within a triangular model, distributional relevance suffices for nonparametric identification of average structural effects via a control function constructed from $F_{X\mid Z}$. We then propose Quantile Least Squares (Q–LS), which aggregates conditional quantiles of $X$ given $Z$ into an optimal mean-square predictor and uses this projection as an instrument in a linear IV estimator. We establish consistency, asymptotic normality, and the validity of standard 2SLS variance formulas, and we discuss regularization across quantiles. Monte Carlo designs show that Q–LS delivers well-centered estimates and near-correct size when mean-based 2SLS suffers from weak instruments. In Health and Retirement Study data, Q–LS exploits Medicare Part D–induced distributional shifts in out-of-pocket risk to sharpen estimates of its effects on depression.
Keywords: Distributional instruments; Quantile least squares; Instrumental variables; Medicare Part D; Out-of-pocket medical spending risk; Depression.
JEL codes: C26; C14; C36; I13; D14.
1 Introduction
A central promise of public health insurance is insurance, not just subsidies: it should protect households against the risk of large, unpredictable medical expenses. In the United States this promise is especially salient for older adults. Nearly all Americans over age 65 are enrolled in Medicare, yet many still face substantial exposure to out-of-pocket (OOP) medical spending, particularly for prescription drugs. [^1] See, for example, Finkelstein and McKnight (2008) on the initial impact of Medicare on OOP spending and Jones et al. (2018) on lifetime medical spending of retirees. The introduction of Medicare Part D in 2006 was explicitly designed to reduce this financial risk by expanding subsidized prescription drug coverage and introducing catastrophic protection. A large empirical literature has since documented how Part D and related coverage changes affect drug utilization, adherence, and OOP spending (e.g. Ketcham and Simon, 2008; Engelhardt and Gruber, 2011; Park and Martin, 2017; Carvalho et al., 2019), and a parallel literature shows that Medicare eligibility and supplemental coverage sharply reduce the probability of catastrophic medical expenditures and financial strain among the elderly (e.g. Barcellos and Jacobson, 2015; Scott et al., 2021; Jones et al., 2019).
Much of this work treats financial exposure itself—for example, the level, variance, or tail probability of OOP spending—as an outcome of policy. From a policy design perspective, however, we often want the reverse object: the causal effect of financial risk exposure on behavior and welfare. Does facing a higher risk of catastrophic prescription drug expenses change adherence, portfolio choice, or retirement decisions? Does improved financial risk protection from Part D or supplemental policies translate into better mental health or reduced financial distress? [^2] See Ayyagari and Shane (2015) and Ayyagari and He (2016) for related evidence on mental health and portfolio choice. Answering such questions requires treating a measure of risk exposure—for example, a catastrophic-spending indicator or a predicted OOP risk index—as an endogenous regressor and using policy variation as an instrument.
In practice, this approach runs into a familiar econometric obstacle. The most natural instruments for a risk index are policy indicators (e.g. Part D eligibility or plan generosity) and features of the plan menu. These instruments primarily shift the distribution of OOP spending—compressing right tails, changing dispersion, and altering the frequency of catastrophic events—while often having relatively modest effects on the mean of the scalar risk measure itself. Empirically, first-stage regressions of a single risk index on these policy variables can have very low $F$-statistics, even when the underlying policy clearly reshapes the full distribution of spending (see, for example, Engelhardt and Gruber, 2011; Barcellos and Jacobson, 2015; Scott et al., 2021).  [^3] Table 9 summarizes some representative studies showing mean effect of insurance on mean OOP spending. Standard IV diagnostics based on mean shifts therefore label the design as “weak” and push researchers toward reduced-form analyses, regression discontinuity designs at age 65, or structural models of plan choice (e.g. Abaluck and Gruber, 2011), rather than causal IV estimates for the effect of financial risk on outcomes.
This paper starts from the observation that, in such settings, the instruments are not weak in any substantive sense: they are strongly relevant for the distribution of the endogenous variable, even if they appear weak for its mean. We formalize this idea by replacing the classical notion of mean relevance—variation in $\mathbb{E}[X\mid Z]$—with a weaker and more general notion of distributional relevance: the conditional distribution $F_{X\mid Z}(\cdot\mid Z)$ of the endogenous variable $X$ shifts with the instrument $Z$ in an $L^{2}$ sense, even when $\mathbb{E}[X\mid Z]$ is constant. In the Medicare example, $X$ may be a scalar index of financial exposure constructed from the OOP distribution, while $Z$ captures policy-induced variation in coverage generosity. Part D and related reforms satisfy distributional relevance whenever they alter the shape (tails, variance, skewness) of the spending distribution, regardless of whether they strongly move its mean.
Our first contribution is to show that distributional relevance is a meaningful notion of instrument strength, not just a descriptive property. Within a nonseparable triangular model, we characterize a class of purely distributional instruments: variables $Z$ for which $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ but $F_{X\mid Z}(\cdot\mid Z)$ is non-degenerate in $L^{2}$. Building on the control-function framework of Chernozhukov et al. (2020), we prove that such instruments still identify average structural effects via the control variable $V=F_{X}(X\mid Z)$, and we relate this notion to existing concepts of instrument strength based on heteroskedasticity (Lewbel, 1997). Thus, even when policy variables are “purely distributional” for a risk index, they can support point identification of its causal effect on outcomes.
Our second, and main, contribution is to develop a practical IV estimator tailored to these designs, which we call Quantile Least Squares (Q–LS). Instead of summarizing the first stage with a conditional mean, Q–LS builds instruments from the entire conditional quantile process of $X$ given $Z$. Concretely, we: (i) estimate a series of first-stage quantile regressions of $X$ on $Z$; (ii) form a data-driven linear combination of these conditional quantiles that best predicts $X$ in mean square error; and (iii) use this optimal quantile-aggregated prediction as the instrument in an otherwise standard linear IV estimator. Under distributional relevance and a mild non-degeneracy condition, the resulting optimal Q–LS instrument is always relevant in the classical sense, even when the conditional mean of $X$ is flat in $Z$. In a “best-case” world where $X\mid Z$ is a pure location shift, we show that Q–LS collapses to optimal 2SLS and attains the same efficiency bound. In more general cases, Q–LS remains reliable precisely when mean-based instruments become weak.
Methodologically, we characterize the asymptotic properties of Q–LS in a linear triangular model with a single endogenous regressor. We show that (i) under distributional relevance and mild regularity conditions, the Q–LS estimator is consistent and asymptotically normal; (ii) standard heteroskedasticity-robust 2SLS standard errors computed with the generated Q–LS instrument are valid for inference; and (iii) ill-posedness in the optimal-weight problem can be handled with ridge or LASSO regularization across quantiles, yielding stable finite-sample performance. We propose simple diagnostics and a decision chart that place Q–LS alongside conventional weak-IV tools: researchers can test for mean relevance, test for distributional relevance, and then choose between classical 2SLS and Q–LS (and between its regularized variants) based on these diagnostics. In this sense, Q–LS is designed to complement, not replace, the existing 2SLS toolkit.
Our Monte Carlo experiments are calibrated to the Medicare context and compare Q–LS to conventional 2SLS and a control-function estimator in designs where instruments primarily shift the distribution of risk rather than its mean. In designs with strong mean relevance, Q–LS and 2SLS deliver nearly identical point estimates and standard errors, confirming that Q–LS does not sacrifice efficiency when classical IV works well. In designs where traditional first-stage $F$-statistics indicate severe weak-instrument problems, Q–LS produces well-centered estimates and near-correct coverage, while 2SLS exhibits substantial bias and undercoverage. Regularized Q–LS variants further stabilise performance when the quantile dictionary is rich relative to sample size.
Finally, we illustrate the usefulness of distributional instruments in the Medicare Part D setting. Using the Health and Retirement Study (HRS), we construct individual-level measures of OOP medical spending risk in real (2015) dollars and exploit Part D as a source of exogenous variation in risk exposure. We show, first, that standard scalar risk measures, such as mean OOP spending or catastrophic-spending indicators, can render an otherwise informative policy design “weak” by classical first-stage diagnostics, even when the upper tail of the OOP distribution clearly shifts. Second, we demonstrate how Q–LS, built from a finite dictionary of conditional quantiles of OOP spending, recovers the same causal effects as linear 2SLS in specifications where the real OOP first stage is strong, while delivering much tighter confidence intervals and more stable estimates in specifications where mean relevance is weak. Substantively, our estimates indicate that higher OOP risk is associated with worse mental health among older Americans. To our knowledge, this is the first empirical illustration of how distributional instruments can materially improve the precision and credibility of IV estimates in applied work on medical spending risk.
Contributions.
Our findings contribute to two main parts of the instrumental‐variables literature. First, on the identification side, building on the control–function results of Chernozhukov et al. (2020), we highlight that their argument continues to apply even when instruments are purely distributional: even when $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ and the conditional mean of the endogenous variable is completely flat in the instrument, nontrivial variation in the conditional distribution $F_{X\mid Z}(\cdot\mid Z)$ is sufficient to identify average structural effects in a nonseparable triangular model via a control function based on $F_{X\mid Z}(X\mid Z)$. Our contribution is to make this “purely distributional” case explicit and to connect it to weak–IV diagnostics in linear models, where instruments may look weak in the mean yet be strong in the distribution. This formalizes a notion of distributional relevance that extends classical mean relevance in a direction that is natural for risk and tail–outcome applications.
Second, on the linear IV side, we develop a simple, implementable estimator, Quantile Least Squares (Q–LS), which constructs an optimal distributional instrument from conditional quantiles and then uses it in a conventional linear IV step. Under a strong-distributional-relevance condition, Q–LS shares the same large-sample behavior as 2SLS when instruments are strong in the mean, but in designs where policy reforms mainly reshape the distribution of risk it can extract a much stronger effective first stage from the same variation, complementing—rather than replacing—existing weak-IV diagnostics and robust inference tools. We characterize its finite–sample behavior through Monte Carlo designs.
Our empirical application illustrates how Q–LS can extract a strong first stage from policy variation that looks weak in classical mean-based diagnostics, while collapsing back to conventional IV when the instrument is strong in the mean.
The remainder of the paper is organized as follows. Section 2 discusses weak identification, quantile regression, and some existing ways to circumvent weak identification. Section 3 introduces the linear IV setup and formalizes distributional relevance. Section 4 defines the Q–LS estimator, presents its asymptotic properties, and compares Q–LS to classical 2SLS and shows that they coincide in a Gaussian/location setting. Section 5 discusses regularization, diagnostics, and practical implementation. Section 6 outlines the simulation designs and results. Section 7 presents the Medicare Part D application. Section 8 concludes.
2 Literature Review
The aim of this research is to develop a general framework that circumvents the classical notion of weak identification of average structural functions in an instrumental variables design. In this section, we first describe and characterize weak identification in triangular models. Second, we summarize quantile regression and discuss how it has been used to address endogeneity in structural functions. Finally, we discuss how our proposed method relates to the existing literature that imposes strong functional form restrictions on the underlying model to circumvent the problem of weak identification.
2.1 Weak identification in triangular and linear IV models
Let $(Y,X,Z)$ be random variables that follow a triangular system
$X=g(Z,\eta),\qquad Y=f(X,\varepsilon),\qquad(\eta,\varepsilon)\perp Z,$
where $f$ and $g$ are unknown measurable structural functions, and $\varepsilon,\eta$ are unobserved disturbances. The vector $Z$ collects the instruments. The average structural function (ASF) is
$\mu(x):=\mathbb{E}\big[f(x,\varepsilon)\big],$
so identification of $\mu(\cdot)$ hinges on how variation in $Z$ propagates to variation in $X$ through $g(\cdot)$ and, ultimately, to variation in $Y$.
In nonseparable triangular models, identification of $\mu(\cdot)$ often proceeds via control-function or distribution-regression arguments that construct a control variable $V=V(X,Z)$ (for example $V=F_{X\mid Z}(X\mid Z)$) such that $Y\perp Z\mid(X,V)$ and suitable completeness conditions hold; see, among others, Newey and Powell (2003); Newey and Stouli (2021), Florens et al. (2008), Imbens and Newey (2009), Chernozhukov et al. (2020), and Dunker et al. (2023). In this framework, weak identification arises when the mapping from $Z$ to the distribution of $X$ is too “flat” to generate informative variation in the control function. Intuitively, if changes in $Z$ induce only very small or nearly collinear shifts in $F_{X\mid Z}(\cdot\mid z)$, then the associated conditional moment conditions are nearly redundant and the inverse problem that recovers $\mu(\cdot)$ becomes ill–posed or nearly unidentified. This is the nonlinear analogue of weak instruments: the conditional distribution of $X$ given $Z$ is only weakly sensitive to $Z$, so the control function has little effective variation.
To connect with the more familiar linear IV setting, consider the linear special case
$Y=X\beta+W^{\prime}\gamma+u,\qquad X=Z^{\prime}\pi+W^{\prime}\delta+v,$
where $Z$ are excluded instruments, $W$ are included exogenous regressors, and $\mathbb{E}[u\mid Z,W]\neq 0$ while $\mathbb{E}[v\mid Z,W]=0$. Classical weak-instrument problems arise when the first-stage coefficients $\pi$ are small, so that the concentration parameter is close to zero and the instruments explain only a tiny fraction of the variation in $X$; see Staiger and Stock (1997), Stock and Yogo (2005), and Dufour (1997) for canonical analyses. In this case, 2SLS and related estimators exhibit large finite-sample bias toward the OLS estimand, heavy-tailed or multimodal sampling distributions, and nonstandard asymptotic behavior ((Andrews and Mikusheva, 2014; Andrews et al., 2019) for a general overview). These issues are amplified in many-instrument settings (e.g. Bekker, 1994; Hansen et al., 2008), where the effective first-stage signal per instrument becomes small even if the joint $F$ is sizable.
The weak-identification phenomena in triangular models are conceptually similar but operate at the level of distributions rather than means. Instead of small first-stage coefficients on $Z$ in a linear projection of $X$ on $Z$, the problem is that the instrument-induced shifts in $F_{X\mid Z}(\cdot\mid z)$ are too limited or nearly fail completeness, so the mapping from structural objects to observables is nearly singular. As emphasized by Lewbel (2019), both linear and nonlinear models admit a continuum of identification strengths, ranging from strong point identification with regular asymptotics, through weak or set identification with nonstandard limits, to complete lack of identification.
Our Q–LS estimator can be viewed as a two-step device that first maps the original instrument vector $Z$ into a scalar distributional instrument
$h(Z)\in\mathcal{H}=\Big\{\int_{0}^{1}\omega(\tau)Q_{X\mid Z}(\tau\mid Z)\,d\tau:\omega\in L^{2}(0,1)\Big\},$
and then applies conventional linear IV using $h(Z)$ in place of $Z$. From this perspective, Assumption 1 is the analogue of a strong-IV condition on the generated instrument $h(Z)$: it rules out sequences of designs in which the covariance between $X$ and $h(Z)$ drifts to zero at a $\sqrt{n}$–local rate.
We therefore do not claim that Q–LS is generally robust in the weak-IV sense of Staiger and Stock (1997) or that it uniformly dominates Anderson–Rubin or conditional tests (e.g. Andrews et al., 2019; Olea and Pflueger, 2013) when instruments are arbitrarily weak. Rather, the contribution of our distributional relevance framework is to highlight a different margin of strength: designs in which $\operatorname{Var}(\mathbb{E}[X\mid Z])$ is small or zero, so that instruments appear weak for the mean, may still be strong for the distribution, with $\mathbb{E}[h_{\mathrm{opt}}(Z)^{2}]$ bounded away from zero. In such cases, Q–LS leverages distributional variation in $X\mid Z$ to construct an effective instrument $h(Z)$ that satisfies a strong-IV condition even when the original mean-based instrument fails conventional first-stage diagnostics.
Our notion of distributional relevance can be viewed as specifying a “strong-instrument” regime for triangular models at the distributional level: it requires that $F_{X\mid Z}(\cdot\mid z)$ varies nontrivially with $Z$ in an $L^{2}$ sense, even when the conditional mean $\mathbb{E}[X\mid Z]$ is flat. This allows us to focus on designs where instruments may be weak for linear first stages but remain informative about the distribution of $X$, and hence can still support identification and efficient estimation once the first stage is recast in distributional terms.
2.2 Quantile regression and endogeneity
Quantile regression was first formalised as an optimization problem minimizing asymmetrically weighted absolute residuals by Koenker and Bassett (1978), to estimate any chosen quantile $\tau\in(0,1)$ of the distribution of an outcome variable $X$ conditional on a set of exogenous regressors $Z$,
$Q_{X|Z}\big(\tau\big|Z=z\big)=\inf\{X:F_{X|Z}(x|z)\geq\tau\},$
where $F_{X|Z}(x|z)$ is the conditional cumulative distribution function (CDF) of $X$ given $Z=z$.
The linear quantile regression model specifies
$Q_{X}\big(\tau\big|Z=z\big)=Z^{\prime}\pi(\tau),$
where $\pi(\tau)$ is estimated by solving the expected check-loss function
$\min_{\pi}\mathbb{E}\big[\rho_{\tau}(X-Z^{\prime}\pi)\big],\;\;\;\rho_{\tau}(u)=u\big(\tau-\mathbbm{1}\{u<0\}\big)$
This objective function is convex but non-differentiable, yielding a linear programming problem that can be solved efficiently in moderately large samples.
Identification of quantile regression relies on mild regularity conditions, including correct specification of the conditional quantile function, sufficient variation in the regressors, and a rank condition ensuring uniqueness of the conditional quantile. These conditions do not require parametric assumptions on the error distribution
Abadie et al. (2002) were the first to combine quantile methods with instrumental variables to estimate distributional treatment effects—that is, how a binary treatment affects quantiles of the outcome distribution—under endogeneity. Chernozhukov and Hansen (2005) generalized this approach to estimate treatment effects at different points of the outcome distribution (quantile treatment effects) under endogeneity and introduced the notion of the structural quantile function. See Wüthrich (2020) for a detailed comparison of Chernozhukov and Hansen (2005) and Abadie et al. (2002). Since these seminal contributions, many variants of quantile instrumental variable methods have been proposed. For example, Powell (2020) develop an instrumental-variable approach to estimate unconditional quantile treatment effects, while Chetverikov et al. (2016) propose methods for group-level quantile treatment effects.
There is also a growing literature that integrates over a grid of quantiles to estimate endogenous structural functions. Chernozhukov et al. (2015) propose a Censored Quantile Instrumental Variables (CQIV) estimator that uses a control-function approach to estimate structural quantiles of a censored outcome with endogenous regressors. In their framework, the control function is estimated by integrating over a grid of quantiles. Chernozhukov et al. (2020) extend the control-function framework of Chernozhukov et al. (2015) and develop a unified approach for estimating average, quantile, and distributional structural functions. The control-function estimator we propose for nonseparable triangular models estimates the control function using the same strategy as Chernozhukov et al. (2020). See Appendix A for further details.
2.3 Existing ways to circumvent weak identification
There is a substantial literature that seeks to circumvent the problem of weak identification by imposing specific functional form restrictions. In this section, we discuss several key strands of this literature and how it relates to our proposed notion of distributional relevance.
There is a growing literature that imposes functional form restrictions on higher moments. Lewbel (1997) show that, in linear models with measurement error and endogeneity, instruments can be constructed from second-moment variation:
$Z_{i}^{\text{Lev}}=(Z_{i}-\bar{Z})(X_{i}-\bar{X}),$
provided that the variation of $X$ conditional on $Z$ changes with $Z$ even when $\mathbb{E}[X\mid Z]$ is constant. This identification strategy exploits variance shifts. Several studies have extended Lewbel (1997) and used second-moment restrictions to achieve identification under weaker functional form restrictions; see, for example, Klein and Vella (2010); Lewbel (2012). These restrictions can be viewed as a specific form of distributional relevance (variance shifts) and are therefore encompassed by our framework.
Tsyawo (2022) also impose a functional form restriction that ensures identification in models of the form $\mathbb{E}[Y-X\beta\mid Z]=0$ without excluded instruments. Specifically, Tsyawo (2022) assume a nonlinear completeness condition requiring injectivity of the conditional mean operator:
$\mathbb{E}[X\mid Z]\tau=0\quad\Rightarrow\quad\tau=0.$
Completeness relies on sufficiently rich nonlinear mean dependence of $X$ conditional on $Z$, that is, on the mapping $X\mapsto\mathbb{E}[X\mid Z]$.
Our approach is fundamentally different from Tsyawo (2022). Identification does not rely on the conditional mean operator or its injectivity. Instead, distributional variation in $X$ suffices to identify the average structural function $\mu(x)$, even when $\mathbb{E}[X\mid Z]$ is constant and completeness fails. Consequently, our results do not require completeness, mean dependence, or any functional-form restrictions of the type imposed by Tsyawo (2022).
Babii and Florens (2025) study estimation in nonidentified linear inverse problems of the from $K\varphi=r$, where $K$ is a compact operator between Hilbert spaces. In nonparametric IV models, $K$ corresponds to the conditional expectation operator $(K\varphi)(z)=\mathbb{E}[\varphi(X)\mid Z=z]$. When completeness (injectivity of $K$) fails, $\varphi$ is not point-identified, but spectral regularization methods converge to the “best approximation” $\varphi_{1}$ in the orthogonal complement of the null space of $K$. Their work characterizes risk bounds and asymptotic distributions under varying degrees of identification.
Our framework differs in two central respects. First, identification is not formulated as an inverse problem with operator $K\varphi=r$. Instead, we use a structural triangular model with distributional relevance, which yields full nonparametric identification of the average structural function $\mu(x)$ when the usual nonparametric IV completeness condition fails. Second, whereas Babii and Florens (2025) obtain convergence to a best approximation under nonidentification, our results deliver point identification of $\mu(x)$ by exploiting the rank structure of the first stage and the distributional variation of $X$. Consequently, our contribution is complementary: we provide an identification and estimation framework that does not rely on operator injectivity or completeness, and remains applicable with instruments that are mean-irrelevant but distributionally strong.
3 Linear IV setup and distributional relevance
3.1 Linear IV model
We consider the linear triangular model
$Y=\alpha+\beta X+Z_{1}^{\prime}\gamma+\varepsilon,\qquad\mathbb{E}[\varepsilon\mid Z]=0,$ (1)
where $Y$ is the outcome, $X$ is a scalar endogenous regressor, $Z_{1}$ is a vector of exogenous covariates included in the outcome equation, and $Z=(Z_{1},Z_{2})$ collects $Z_{1}$ and a vector of excluded instruments $Z_{2}$.
In the classical IV framework, instrument strength is measured by variation in the conditional mean of $X$ given $Z$:
$m(Z):=\mathbb{E}[X\mid Z],\qquad\operatorname{Var}(m(Z))>0.$
When $\operatorname{Var}(m(Z))$ is close to zero, instruments are labelled “weak” and standard 2SLS inference is unreliable.
Our aim is to replace this mean-based notion of strength with a weaker and more general concept based on the full conditional distribution of $X$.
3.2 Distributional relevance
Let
$F_{X\mid Z}(x\mid Z):=\mathbb{P}(X\leq x\mid Z),\qquad\bar{F}_{X}(x):=\mathbb{E}\big[F_{X\mid Z}(x\mid Z)\big]$
denote the conditional and unconditional distribution functions of $X$. We measure variation in $F_{X\mid Z}(\cdot\mid Z)$ using an $L^{2}$ norm over $x$. For concreteness, let $\lambda$ be Lebesgue measure on a compact interval containing the support of $X$, and define
$\|F_{X\mid Z}(\cdot\mid Z)\|_{L^{2}(\lambda)}^{2}:=\int\big(F_{X\mid Z}(x\mid Z)\big)^{2}\,d\lambda(x).$
Definition 1 (Mean relevance).
We say that the instrument $Z$ is mean-relevant for $X$ if
$\operatorname{Var}\big(\mathbb{E}[X\mid Z]\big)=\mathbb{E}\Big[\big(\mathbb{E}[X\mid Z]-\mathbb{E}[X]\big)^{2}\Big]>0.$
Definition 2 (Distributional relevance).
We say that the instrument $Z$ is distributionally relevant for $X$ if
$\mathbb{E}\Big[\big\|F_{X\mid Z}(\cdot\mid Z)-\bar{F}_{X}(\cdot)\big\|_{L^{2}(\lambda)}^{2}\Big]>0.$ (2)
Equivalently, there exist $z_{p},z_{q}$ in the support of $Z$ such that
$\big\|F_{X\mid Z}(\cdot\mid z_{p})-F_{X\mid Z}(\cdot\mid z_{q})\big\|_{L^{2}(\lambda)}>0.$
Any change in $\mathbb{E}[X\mid Z]$ across $Z$ induces a difference in $F_{X\mid Z}(\cdot\mid Z)$, so mean relevance implies distributional relevance. The converse need not hold: $Z$ can be distributionally relevant even when $\mathbb{E}[X\mid Z]$ is constant in $Z$. This motivates the following concept.
Definition 3 (Purely distributional instruments).
We say that $Z$ is a purely distributional instrument for $X$ if
  1. (i)
    $Z$ is distributionally relevant for $X$ in the sense of Definition 2, and
  2. (ii)
    $Z$ is mean-irrelevant for $X$, i.e.
    $\operatorname{Var}(\mathbb{E}[X\mid Z])=0\quad\Longleftrightarrow\quad\mathbb{E}[X\mid Z=z]=\mathbb{E}[X]\quad\text{for all }z.$
Classical IV analysis treats mean relevance—variation in $\mathbb{E}[X\mid Z]$—as the benchmark notion of instrument strength. Distributional relevance weakens this requirement: it allows $\mathbb{E}[X\mid Z]$ to be completely flat in $Z$, as long as the conditional distribution $F_{X\mid Z}(\cdot\mid Z)$ varies nontrivially in $L^{2}$. In this sense, an instrument can be “strong” for the distribution of $X$ even when it is “weak” for its mean. We refer to instruments that satisfy distributional relevance but have $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ as purely distributional: they convey no information about $\mathbb{E}[X\mid Z]$ but still transmit rich information about the risk environment through higher moments and tail behavior.
In the linear IV analysis that follows, distributional relevance and the linear quantile representation for $X\mid Z$ guarantee the existence of an optimal Q–LS instrument that is relevant in the classical sense unless the projection degenerates. In the nonseparable triangular model in Appendix A, the condition that $Z$ is purely distributional is enough for the identification of the average structural function via a control function based on $F_{X\mid Z}(X\mid Z)$.
Remark 1 (Illustrative example).
Let $Z_{2}\in\{0,1\}$ be a binary instrument with $\mathbb{P}(Z_{2}=0)=\mathbb{P}(Z_{2}=1)=1/2$, and let $U\sim\mathcal{N}(0,1)$ be independent of $Z_{2}$. Define
$X=\sigma(Z_{2})\,U,\qquad\sigma(0)=1,\ \sigma(1)=2.$
Then $\mathbb{E}[X\mid Z_{2}=z_{2}]=0$ for all $z_{2}$, so $\operatorname{Var}(\mathbb{E}[X\mid Z_{2}])=0$ and $Z_{2}$ is mean-irrelevant. However, the conditional distributions differ: $X\mid Z_{2}=0\sim\mathcal{N}(0,1)$ and $X\mid Z_{2}=1\sim\mathcal{N}(0,4)$, so $F_{X\mid Z_{2}}(\cdot\mid 0)\neq F_{X\mid Z_{2}}(\cdot\mid 1)$ in $L^{2}$. Thus $Z_{2}$ is a purely distributional instrument. In a classical first-stage regression of $X$ on $Z_{2}$, the coefficient on $Z_{2}$ would be zero, but the instrument carries substantial information about the dispersion of $X$.
In what follows, distributional relevance will be our primitive assumption on the instrument. We show that, under a linear quantile representation for $X\mid Z$, it is enough to construct a strong instrument as a linear functional of the conditional quantile process, and to obtain consistent and asymptotically normal estimates of $(\alpha,\beta,\gamma)$ in the linear model (1).
4 The Q–LS estimator in the linear model
We now define the Quantile Least Squares IV (Q–LS) estimator. The key object is an optimal instrument constructed as a linear functional of the conditional quantile process of $X\mid Z$.
4.1 Conditional quantiles and quantile-aggregated instruments
We work with the conditional quantile function of $X$ given $Z$. For each $\tau\in(0,1)$, assume
$Q_{X\mid Z}(\tau\mid Z)=Z^{\prime}\pi_{0}(\tau),\qquad\tau\in(0,1),$ (3)
where $\pi_{0}(\tau)\in\mathbb{R}^{\dim(Z)}$ is a measurable function of $\tau$. We write
$g(\tau,Z):=Q_{X\mid Z}(\tau\mid Z)=Z^{\prime}\pi_{0}(\tau).$
Under mild regularity, distributional relevance is equivalent to the statement that the conditional quantile process $\{g(\tau,Z):\tau\in(0,1)\}$ is non-constant in $Z$ in $L^{2}$. [^4] For concreteness, we implement Q–LS using linear quantile regressions of $X$ on $Z$ at a finite grid of quantile indices. This choice is purely for illustration: any alternative specification for the conditional quantiles (e.g., nonlinear, spline or series-based, or machine–learning first stages) could be used to construct the dictionary $G(Z)$ and hence the Q–LS instrument.
We restrict attention to instruments that are linear functionals of this quantile process. Let $\omega(\cdot)$ be a square-integrable weight function on $(0,1)$. We define the quantile-aggregated instrument
$h_{\omega}(Z):=\int_{0}^{1}\omega(\tau)\,g(\tau,Z)\,d\tau=\int_{0}^{1}\omega(\tau)\,Q_{X\mid Z}(\tau\mid Z)\,d\tau.$ (4)
The corresponding moment conditions for the structural parameter $\theta:=(\alpha,\beta,\gamma^{\prime})^{\prime}$ are
$\mathbb{E}\!\left[\begin{pmatrix}h_{\omega}(Z)\\ Z_{1}\end{pmatrix}\big(Y-\alpha-\beta X-Z_{1}^{\prime}\gamma\big)\right]=0.$ (5)
Let
$\mathcal{H}:=\Big\{h_{\omega}(\cdot):h_{\omega}(Z)=\int_{0}^{1}\omega(\tau)\,g(\tau,Z)\,d\tau,\ \omega\in L^{2}(0,1)\Big\}$
be the linear span (in $L^{2}$) of the conditional quantile process.
4.2 Optimal Q–LS and relevance under distributional relevance
Following the classical optimal IV logic for a single endogenous regressor, we consider the element of $\mathcal{H}$ that best predicts $X$ in mean square error.
Definition 4 (Optimal Q–LS).
The optimal Q–LS weight function $\omega_{\mathrm{opt}}(\cdot)$ is any solution to
$\omega_{\mathrm{opt}}\in\arg\min_{\omega\in L^{2}(0,1)}\mathbb{E}\!\left[\big(X-h_{\omega}(Z)\big)^{2}\right],$ (6)
and the associated optimal Q–LS instrument is
$h_{\mathrm{opt}}(Z):=h_{\omega_{\mathrm{opt}}}(Z)=\int_{0}^{1}\omega_{\mathrm{opt}}(\tau)\,Q_{X\mid Z}(\tau\mid Z)\,d\tau.$ (7)
By construction, $h_{\mathrm{opt}}(Z)$ is the $L^{2}$-projection of $X$ onto the closed linear span $\mathcal{H}$ of quantile-generated instruments. The next lemma shows that, whenever this projection is nonzero, the resulting instrument is automatically relevant for $X$ in the classical sense.
Lemma 1 (Distributional relevance and optimal Q–LS).
Suppose Assumption 2 and (3) hold, and let $h_{\mathrm{opt}}$ be defined as in Definition 4. Then
$X=h_{\mathrm{opt}}(Z)+r(Z),\qquad r\perp\mathcal{H},$
and either
$h_{\mathrm{opt}}(Z)=0\quad\text{a.s.},$
or
$\operatorname{cov}\big(h_{\mathrm{opt}}(Z),X\big)=\mathbb{E}\big[h_{\mathrm{opt}}(Z)^{2}\big]>0.$
In particular, whenever $h_{\mathrm{opt}}$ is non-degenerate it is a relevant instrument for $X$.
Proof.
By definition of $h_{\mathrm{opt}}$ as the $L^{2}$-projection of $X$ onto the closed linear subspace $\mathcal{H}$, we can write
$X=h_{\mathrm{opt}}(Z)+r(Z),$
where $r(Z)$ is the projection residual and satisfies $\mathbb{E}[h(Z)\,r(Z)]=0$ for all $h\in\mathcal{H}$. In particular, $\mathbb{E}[h_{\mathrm{opt}}(Z)\,r(Z)]=0$, so
$\mathbb{E}[X\,h_{\mathrm{opt}}(Z)]=\mathbb{E}\big[(h_{\mathrm{opt}}(Z)+r(Z))\,h_{\mathrm{opt}}(Z)\big]=\mathbb{E}\big[h_{\mathrm{opt}}(Z)^{2}\big]+\mathbb{E}\big[h_{\mathrm{opt}}(Z)\,r(Z)\big]=\mathbb{E}\big[h_{\mathrm{opt}}(Z)^{2}\big].$
Thus either $h_{\mathrm{opt}}(Z)=0$ a.s., in which case both sides vanish, or $\mathbb{E}[h_{\mathrm{opt}}(Z)^{2}]>0$ and $h_{\mathrm{opt}}$ is correlated with $X$. ∎
Lemma 1 delivers a clean dichotomy: the optimal Q–LS instrument is either identically zero or, if nonzero, it is automatically relevant and its “strength” is measured by $\mathbb{E}[h_{\mathrm{opt}}(Z)^{2}]$. Distributional relevance guarantees that $\mathcal{H}$ contains more than just constants, but it does not by itself rule out the degenerate case $h_{\mathrm{opt}}\equiv 0$. A simple example is a pure variance-shift design,
$X=\sigma(Z)\,U,\qquad\mathbb{E}[U]=0,\quad U\perp\!\!\!\perp Z,$
with $\sigma(Z)$ non-constant. In this case the conditional distributions $F_{X\mid Z}(\cdot\mid Z)$ vary with $Z$ (distributional relevance), yet $\operatorname{cov}(X,h(Z))=0$ for every $h\in\mathcal{H}$ so that $h_{\mathrm{opt}}(Z)\equiv 0$. Our Design B in Section 6 provides a Monte Carlo illustration of this phenomenon.
For identification, we therefore add an explicit relevance condition that is the exact analogue of the usual mean-relevance condition in classical IV.
Assumption 1 (Strong distributional relevance).
Under the linear quantile representation (3), the optimal Q–LS instrument $h_{\mathrm{opt}}(Z)$ defined in (7) satisfies
$\mathbb{E}\big[h_{\mathrm{opt}}(Z)^{2}\big]\geq c>0$
for some constant $c$ that does not depend on the sample size $n$.
Assumption 1 is the distributional analogue of the usual “strong IV” condition in linear models, where one assumes that the first-stage coefficient on the excluded instruments does not drift to zero as $n$ grows (e.g. Staiger and Stock, 1997; Stock and Yogo, 2005. Here we impose a lower bound on the variance of the optimal Q–LS instrument $h_{\mathrm{opt}}(Z)$, ruling out sequences of designs in which instruments become weak either in the mean or in the distribution. Throughout our asymptotic analysis we work under Assumption 1, so our results should be interpreted as strong-distributional-IV asymptotics.
Under Assumption 1, Lemma 1 implies that $h_{\mathrm{opt}}$ is a non-degenerate, relevant instrument constructed purely from the conditional distribution of $X\mid Z$, even when $\mathbb{E}[X\mid Z]$ is constant in $Z$.
Throughout the linear IV analysis, we assume that distributional relevance holds and that $X\mid Z$ admits the linear quantile representation in (3). Under these conditions, Lemma 1 shows that the optimal Q–LS instrument $h_{\mathrm{opt}}(Z)$ is either identically zero or a classically relevant instrument for $X$, and Assumption 1 rules out the degenerate case.
4.3 Finite-grid implementation
In practice, we approximate the integral in (4) using a finite grid of quantiles. Let
$0<\tau_{1}<\cdots<\tau_{K}<1$
be a set of quantile indexes with gap of order $1/K$, and let $\Delta\tau_{k}$ denote the associated quadrature weights (for simplicity, one can take $\Delta\tau_{k}=1/K$).
For each $\tau_{k}$, we estimate the first-stage quantile regression
$\widehat{\pi}_{n}(\tau_{k})\in\arg\min_{\pi}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau_{k}}\big(X_{i}-Z_{i}^{\prime}\pi\big),\qquad k=1,\dots,K,$ (8)
and define the estimated quantile-based instruments
$\widehat{g}_{k}(Z_{i}):=Z_{i}^{\prime}\widehat{\pi}_{n}(\tau_{k}),\qquad i=1,\dots,n,\quad k=1,\dots,K.$
Stacking them, let
$\widehat{G}_{i}:=\big(\widehat{g}_{1}(Z_{i}),\dots,\widehat{g}_{K}(Z_{i})\big)^{\prime},\qquad\widehat{G}:=\begin{pmatrix}\widehat{G}_{1}^{\prime}\\ \vdots\\ \widehat{G}_{n}^{\prime}\end{pmatrix}\in\mathbb{R}^{n\times K}.$
Discrete Q–LS weights.
A natural discrete analogue of (6) chooses weights $w\in\mathbb{R}^{K}$ to best predict $X$ from the $K$ quantile instruments:
$\widehat{w}_{n}\in\arg\min_{w\in\mathbb{R}^{K}}\frac{1}{n}\sum_{i=1}^{n}\big(X_{i}-\widehat{G}_{i}^{\prime}w\big)^{2},$ (9)
so that, when $\widehat{G}^{\prime}\widehat{G}$ is invertible,
$\widehat{w}_{n}=(\widehat{G}^{\prime}\widehat{G})^{-1}\widehat{G}^{\prime}X,\qquad X=(X_{1},\dots,X_{n})^{\prime}.$ (10)
The corresponding finite-sample Q–LS first-stage prediction is
$\widehat{X}_{i}^{\mathrm{Q\text{--}LS}}:=\widehat{G}_{i}^{\prime}\widehat{w}_{n}=\sum_{k=1}^{K}\widehat{w}_{n,k}\,\widehat{g}_{k}(Z_{i}),\qquad i=1,\dots,n.$ (11)
This $\widehat{X}_{i}^{\mathrm{Q\text{--}LS}}$ is the empirical best linear predictor of $X_{i}$ in the span of the $K$ quantile-based instruments.
Second stage.
Let $y=(Y_{1},\dots,Y_{n})^{\prime}$ be the outcome vector, and define the regressor and instrument matrices
$S=[\mathbf{1},\,X,\,Z_{1}],\qquad M=[\widehat{X}^{\mathrm{Q\text{--}LS}},\,Z_{1}],$
where $\widehat{X}^{\mathrm{Q\text{--}LS}}=(\widehat{X}_{1}^{\mathrm{Q\text{--}LS}},\dots,\widehat{X}_{n}^{\mathrm{Q\text{--}LS}})^{\prime}$ and $\mathbf{1}$ denotes the $n$-vector of ones. The Q–LS estimator of $\theta=(\alpha,\beta,\gamma^{\prime})^{\prime}$ is the usual 2SLS estimator
$\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}:=\big(S^{\prime}P_{M}S\big)^{-1}S^{\prime}P_{M}y,\qquad P_{M}=M(M^{\prime}M)^{-1}M^{\prime}.$ (12)
4.4 Asymptotic properties
We now state conditions under which the Q–LS estimator is consistent and asymptotically normal. The key identification condition is distributional relevance, which, via Lemma 1, implies relevance of the optimal Q–LS instrument.
Assumption 2 (Regularity and identification).
  1. (i)
    The sample $\{(Y_{i},X_{i},Z_{i})\}_{i=1}^{n}$ is i.i.d., and $\mathbb{E}\|Z\|^{2}<\infty$, $\mathbb{E}|X|^{2}<\infty$, $\mathbb{E}|Y|^{2}<\infty$.
  2. (ii)
    For each $\tau\in(0,1)$, the conditional quantile $Q_{X\mid Z}(\tau\mid Z)$ exists and satisfies (3), with $\pi_{0}(\cdot)$ continuous on $(0,1)$ and $\int_{0}^{1}\|\pi_{0}(\tau)\|^{2}d\tau<\infty$.
  3. (iii)
    The structural error satisfies $\mathbb{E}[\varepsilon\mid Z]=0$ and $\mathbb{E}[\varepsilon^{2}\mid Z]<\infty$.
  4. (iv)
    Distributional relevance holds in the sense of Definition 2, and the span $\mathcal{H}$ generated by $\{g(\tau,Z):\tau\in(0,1)\}$ is not reduced to constants. Let $h_{\mathrm{opt}}$ be defined as in Definition 4, and suppose the population matrix
    $\mathbb{E}\big[S_{0}(Z)^{\prime}S_{0}(Z)\big],\qquad S_{0}(Z)=(1,X,Z_{1}^{\prime})$
    is nonsingular.
  5. (v)
    The quantile grid $\{\tau_{k}\}_{k=1}^{K}$ is dense in $(0,1)$ as $K=K_{n}\to\infty$, and the number of quantiles satisfies $K_{n}^{2}/n\to 0$ as $n\to\infty$.
Proposition 1 (Consistency of the Q–LS estimator).
Suppose Assumptions 1 and 2 hold, and the weights $\widehat{w}_{n}$ are defined by (9) with $K=K_{n}\to\infty$ and $K_{n}^{2}/n\to 0$. Then
$\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}\xrightarrow{p}\theta\qquad\text{as }n\to\infty.$
Proof.
By Assumption 2(ii) and standard quantile regression theory, we have, for each fixed $\tau$, $\widehat{\pi}_{n}(\tau)\xrightarrow{p}\pi_{0}(\tau)$, and, since the grid $\{\tau_{k}\}$ becomes dense and $K_{n}^{2}/n\to 0$, the array $\{\widehat{\pi}_{n}(\tau_{k})\}$ uniformly approximates $\{\pi_{0}(\tau_{k})\}$ in mean square. It follows that
$\max_{1\leq k\leq K_{n}}\mathbb{E}\big[\big(\widehat{g}_{k}(Z)-g(\tau_{k},Z)\big)^{2}\big]\to 0,$
and hence $\widehat{G}^{\prime}\widehat{G}/n$ converges in probability to the corresponding population limit, and similarly for $\widehat{G}^{\prime}X/n$. Under $K_{n}^{2}/n\to 0$, the estimation error from the first-stage quantile regressions is $o_{p}(n^{-1/2})$ in the moment conditions.
Therefore the discrete weights $\widehat{w}_{n}$ converge in probability to the population weights $w_{0}$ that minimize $\mathbb{E}[(X-G_{0}(Z)^{\prime}w)^{2}]$ in the corresponding finite-dimensional approximation space, where $G_{0}(Z)$ stacks the $g(\tau_{k},Z)$’s. As $K_{n}\to\infty$ and the grid becomes dense, the discrete approximation $G_{0}(Z)^{\prime}w_{0}$ converges in $L^{2}$ to the continuum optimal instrument $h_{\mathrm{opt}}(Z)$ defined in Definition 4. In particular,
$\mathbb{E}\big[\big(\widehat{X}_{i}^{\mathrm{Q\text{--}LS}}-h_{\mathrm{opt}}(Z_{i})\big)^{2}\big]\to 0.$
Standard arguments for 2SLS with generated instruments then imply that $\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}$ converges in probability to the unique solution of the population IV moment condition with instrument $h_{\mathrm{opt}}(Z)$, which is $\theta_{0}$ by Assumption 2(iv). ∎
For asymptotic normality, let
$S_{i}:=\big(1,X_{i},Z_{1i}^{\prime}\big)^{\prime},\qquad M_{i}:=\big(\widehat{X}_{i}^{\mathrm{Q\text{--}LS}},Z_{1i}^{\prime}\big)^{\prime}$
denote the stacked regressor and instrument vectors used in the second stage, and let
$\widehat{\varepsilon}_{i}:=Y_{i}-S_{i}^{\prime}\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}.$
Define the sample matrices
$\widehat{A}_{n}:=\frac{1}{n}\sum_{i=1}^{n}S_{i}M_{i}^{\prime},\qquad\widehat{Q}_{n}:=\frac{1}{n}\sum_{i=1}^{n}M_{i}M_{i}^{\prime},$
and the heteroskedasticity-robust “meat” matrix
$\widehat{\Omega}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\widehat{\varepsilon}_{i}^{\,2}\,M_{i}M_{i}^{\prime}.$
Assumption 3 (Additional conditions for asymptotic normality).
In addition to Assumption 2, suppose that:
  1. (i)
    The fourth moments of $(Y,X,Z)$ are finite: $\mathbb{E}\|Z\|^{4}<\infty$, $\mathbb{E}|X|^{4}<\infty$, $\mathbb{E}|Y|^{4}<\infty$.
  2. (ii)
    The conditional fourth moment of the structural error is finite: $\mathbb{E}[\varepsilon^{4}\mid Z]<\infty$ a.s.
  3. (iii)
    The eigenvalues of $\mathbb{E}[M_{i}^{0}M_{i}^{0\prime}]$ are bounded away from zero and infinity, where $M_{i}^{0}=(h_{\mathrm{opt}}(Z_{i}),Z_{1i}^{\prime})^{\prime}$ is the population instrument vector associated with the optimal Q–LS, and the same holds for $\mathbb{E}[S_{i}S_{i}^{\prime}]$.
Let
$A:=\mathbb{E}[S_{i}M_{i}^{0\prime}],\qquad Q:=\mathbb{E}[M_{i}^{0}M_{i}^{0\prime}],\qquad\Omega:=\mathbb{E}\big[\varepsilon_{i}^{2}M_{i}^{0}M_{i}^{0\prime}\big].$
The population asymptotic covariance matrix of the Q–LS estimator can then be written in the usual IV “sandwich” form
$V_{\mathrm{Q\text{--}LS}}:=\big(AQ^{-1}A^{\prime}\big)^{-1}\big(AQ^{-1}\Omega Q^{-1}A^{\prime}\big)\big(AQ^{-1}A^{\prime}\big)^{-1}.$ (13)
Proposition 2 (Asymptotic normality of the Q–LS estimator).
Suppose Assumptions 1, 2 and 3 hold. Then
$\sqrt{n}\big(\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}-\theta\big)\ \xrightarrow{d}\ \mathcal{N}\big(0,V_{\mathrm{Q\text{--}LS}}\big),$
where $V_{\mathrm{Q\text{--}LS}}$ is given in (13). Moreover, the heteroskedasticity-robust 2SLS covariance matrix computed with the generated Q–LS instrument,
$\widehat{V}_{n}^{\mathrm{Q\text{--}LS}}:=\big(\widehat{A}_{n}\widehat{Q}_{n}^{-1}\widehat{A}_{n}^{\prime}\big)^{-1}\big(\widehat{A}_{n}\widehat{Q}_{n}^{-1}\widehat{\Omega}_{n}\widehat{Q}_{n}^{-1}\widehat{A}_{n}^{\prime}\big)\big(\widehat{A}_{n}\widehat{Q}_{n}^{-1}\widehat{A}_{n}^{\prime}\big)^{-1},$ (14)
is consistent for $V_{\mathrm{Q\text{--}LS}}$, in the sense that $\widehat{V}_{n}^{\mathrm{Q\text{--}LS}}\xrightarrow{p}V_{\mathrm{Q\text{--}LS}}$. Consequently, for any fixed vector $c\in\mathbb{R}^{\dim(\theta_{0})}$,
$\frac{\sqrt{n}\,c^{\prime}\big(\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}-\theta\big)}{\sqrt{c^{\prime}\widehat{V}_{n}^{\mathrm{Q\text{--}LS}}c}}\ \xrightarrow{d}\ \mathcal{N}(0,1).$
Proof.
By Proposition 1, the generated instrument $\widehat{X}_{i}^{\mathrm{Q\text{--}LS}}$ converges in $L^{2}$ to the population optimal instrument $h_{\mathrm{opt}}(Z_{i})$, and the Q–LS estimator solves the sample moment condition
$\frac{1}{n}\sum_{i=1}^{n}M_{i}\big(Y_{i}-S_{i}^{\prime}\theta\big)=0,$
with $M_{i}=(\widehat{X}_{i}^{\mathrm{Q\text{--}LS}},Z_{1i}^{\prime})^{\prime}$. The estimation error in $\widehat{X}_{i}^{\mathrm{Q\text{--}LS}}$ is $o_{p}(n^{-1/2})$ in the associated moment conditions under $K_{n}^{2}/n\to 0$, so the first-order asymptotics coincide with those of a conventional 2SLS estimator that uses the population instrument $M_{i}^{0}=(h_{\mathrm{opt}}(Z_{i}),Z_{1i}^{\prime})^{\prime}$. Standard arguments for IV/GMM with i.i.d. observations, finite fourth moments, and nonsingular Jacobian then yield the stated result, and the consistency of $\widehat{V}_{n}^{\mathrm{Q\text{--}LS}}$ follows by the law of large numbers and Slutsky’s theorem. ∎
In applications, we report standard errors as the square roots of the diagonal elements of $\widehat{V}_{n}^{\mathrm{Q\text{--}LS}}$ in (14), which coincide with the usual heteroskedasticity-robust 2SLS standard errors computed using the generated Q–LS instrument and the exogenous regressors $Z_{1}$ as instruments.
Proposition 1, is derived under Assumption 1, which enforces a strong form of distributional relevance for the generated instrument $h(Z)$. As in the classical linear-IV literature, this rules out weak-IV sequences in which the covariance between the endogenous regressor and the instrument shrinks at a $\sqrt{n}$–local rate. In particular, our asymptotic normality result for Q–LS should be interpreted as a strong-identification limit theory: it guarantees standard Gaussian inference when the optimal distributional instrument $h_{\mathrm{opt}}(Z)$ has nonvanishing variance, but it does not characterize the behavior of Q–LS under general weak-IV sequences. Extending weak-identification robust inference to settings with distributional instruments would require combining our construction with the robust testing approaches of, for example, Andrews et al. (2019) or Olea and Pflueger (2013), which we leave for future work.
4.5 Relation to classical 2SLS
In this section we compare the quantile least squares (Q–LS) estimator to classical two-stage least squares (2SLS) in settings where both procedures are valid.
4.5.1 2SLS and optimal mean-based instruments
In the classical IV framework with a single endogenous regressor, relevance is formulated in terms of the conditional mean
$m(Z):=\mathbb{E}[X\mid Z].$
Under homoskedasticity, the “best” instrument for a single endogenous regressor is the efficient instrument $h^{*}(Z)$, which is proportional to the best $L^{2}$ predictor of $X$ given $Z$, i.e. to $m(Z)$. The optimal 2SLS estimator uses instruments in the linear span of $\{1,Z_{1}^{\prime},m(Z)\}$ and is semiparametrically efficient for $\theta_{0}$ within this mean-based class.
4.5.2 Q–LS versus 2SLS in a location model
By contrast, Q–LS is built from the entire conditional distribution of $X\mid Z$ via its conditional quantiles. Under (3), we construct quantile-aggregated instruments of the form
$h_{\omega}(Z)=\int_{0}^{1}\omega(\tau)\,Q_{X\mid Z}(\tau\mid Z)\,d\tau,$
and define the optimal Q–LS instrument $h_{\mathrm{opt}}$ as in Definition 4, i.e. as the best $L^{2}$ predictor of $X$ within the quantile-generated class $\mathcal{H}$. The Q–LS estimator $\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}$ is the 2SLS estimator that uses $(h_{\mathrm{opt}}(Z),Z_{1})$ as instruments.
The following proposition shows that, in a Gaussian/location setting with mean relevance, Q–LS and optimal 2SLS are asymptotically equivalent.
Proposition 3 (Q–LS versus 2SLS in a location model).
Suppose that, in addition to Assumptions 23, the conditional distribution of $X\mid Z$ belongs to a location family, i.e. there exists a random variable $U$ with $\mathbb{E}[U]=0$ such that
$X\mid Z\ \overset{d}{=}\ m(Z)+U,$
and the distribution of $U$ does not depend on $Z$. Then:
  1. (i)
    For each $\tau\in(0,1)$ we have
    $Q_{X\mid Z}(\tau\mid Z)=m(Z)+c(\tau),$
    for some scalar function $c(\cdot)$ independent of $Z$.
  2. (ii)
    Any quantile-aggregated instrument $h_{\omega}(Z)$ can be written as
    $h_{\omega}(Z)=a_{\omega}\,m(Z)+b_{\omega},\qquad a_{\omega}=\int_{0}^{1}\omega(\tau)\,d\tau,\ \ b_{\omega}=\int_{0}^{1}\omega(\tau)\,c(\tau)\,d\tau.$
  3. (iii)
    The optimal Q–LS instrument $h_{\mathrm{opt}}(Z)$ is proportional to $m(Z)$ and, up to an irrelevant rescaling of the instrument, the Q–LS estimator coincides with the optimal 2SLS estimator:
    $\widehat{\theta}_{n}^{\mathrm{Q\text{--}LS}}\ \equiv\ \widehat{\theta}_{n}^{\mathrm{2SLS}}+o_{p}(n^{-1/2}),$
    so that $V_{\mathrm{Q\text{--}LS}}=V_{\mathrm{2SLS}}$.
Proof.
Under the location assumption, $X\mid Z$ has distribution equal to $U$ shifted by $m(Z)$, so its $\tau$-quantile satisfies $Q_{X\mid Z}(\tau\mid Z)=m(Z)+c(\tau)$, where $c(\tau)$ is the $\tau$-quantile of $U$ and does not depend on $Z$. Any $h_{\omega}\in\mathcal{H}$ is therefore of the form $a_{\omega}m(Z)+b_{\omega}$ as stated. Minimizing $\mathbb{E}[(X-h_{\omega}(Z))^{2}]$ over $\omega$ then reduces to minimizing $\mathbb{E}[(X-am(Z)-b)^{2}]$ over $(a,b)$, whose unique solution is $a=1$ and $b=0$ (up to the usual scale normalization of instruments), so $h_{\mathrm{opt}}(Z)$ is proportional to $m(Z)$. The Q–LS estimator thus corresponds to 2SLS with the optimal mean-based instrument and has the same asymptotic distribution. ∎
Proposition 3 shows that, in a conventional Gaussian or location setting with mean relevance, Q–LS and optimal 2SLS are asymptotically equivalent: they use instruments that differ only by an irrelevant scale factor and attain the same efficiency bound.
Lemma 2 (Equivalence between 2SLS and plug-in OLS on the projected regressor).
Let $\{(Y_{i},X_{i},Z_{1i},Z_{2i})\}_{i=1}^{n}$ be a sample and consider the linear structural model
$Y_{i}=\alpha+\beta X_{i}+Z_{1i}^{\prime}\gamma+\varepsilon_{i},\qquad\mathbb{E}[\varepsilon_{i}\mid Z_{i}]=0,$
with $Z_{i}=(Z_{1i}^{\prime},Z_{2i}^{\prime})^{\prime}$. Let
$S:=\begin{bmatrix}\mathbf{1}&X&Z_{1}\end{bmatrix}\in\mathbb{R}^{n\times(1+1+d_{1})},\qquad M:=\begin{bmatrix}Z_{1}&G\end{bmatrix}\in\mathbb{R}^{n\times(d_{1}+K)},$
where $G$ collects a set of $K$ instrument functions (e.g. quantile-based instruments) and $Z_{1}$ denotes the included exogenous regressors. Let
$P_{M}:=M(M^{\prime}M)^{-1}M^{\prime}$
be the projection matrix onto the column space of $M$ and define the projected regressor
$\tilde{X}:=P_{M}X.$
Consider the following two estimators for $\theta=(\alpha,\beta,\gamma^{\prime})^{\prime}$:
  1. (i)
    The 2SLS estimator with instrument matrix $M$:
    $\widehat{\theta}^{\mathrm{2SLS}}:=(S^{\prime}P_{M}S)^{-1}S^{\prime}P_{M}Y.$
  2. (ii)
    The plug-in OLS estimator based on the projected regressor $\tilde{X}$:
    $\widehat{\theta}^{\mathrm{OLS-proj}}:=\big(\tilde{S}^{\prime}\tilde{S}\big)^{-1}\tilde{S}^{\prime}Y,\qquad\tilde{S}:=\begin{bmatrix}\mathbf{1}&\tilde{X}&Z_{1}\end{bmatrix}$.
Then, in finite samples,
$\widehat{\theta}^{\mathrm{2SLS}}=\widehat{\theta}^{\mathrm{OLS-proj}}.$
In particular, the 2SLS coefficient on $X$ coincides with the OLS coefficient on $\tilde{X}$ in the regression of $Y$ on $(\tilde{X},Z_{1})$.
Proof.
By definition of $P_{M}$, we have
$\tilde{S}:=P_{M}S=\begin{bmatrix}P_{M}\mathbf{1}&P_{M}X&P_{M}Z_{1}\end{bmatrix}$.
Because each column of $Z_{1}$ is in the column space of $M$, $P_{M}Z_{1}=Z_{1}$. Similarly, if the constant is included in $Z_{1}$ (or in $M$), then $P_{M}\mathbf{1}=\mathbf{1}$. By definition, $P_{M}X=\tilde{X}$. Hence $\tilde{S}=[\mathbf{1},\,\tilde{X},\,Z_{1}]$ as defined above.
The 2SLS estimator can be written as
$\widehat{\theta}^{\mathrm{2SLS}}=(S^{\prime}P_{M}S)^{-1}S^{\prime}P_{M}Y.$
Using $\tilde{S}=P_{M}S$, we have
$S^{\prime}P_{M}S=S^{\prime}\tilde{S}=\tilde{S}^{\prime}\tilde{S},\qquad S^{\prime}P_{M}Y=S^{\prime}\tilde{Y}=\tilde{S}^{\prime}Y,$
since $\tilde{Y}:=P_{M}Y$ and $P_{M}$ is symmetric and idempotent. Therefore
$\widehat{\theta}^{\mathrm{2SLS}}=(\tilde{S}^{\prime}\tilde{S})^{-1}\tilde{S}^{\prime}Y=\widehat{\theta}^{\mathrm{OLS-proj}},$
which proves the claim. ∎
Remark 2 (Application to Q–LS).
In the Q–LS setting, let $G$ collect the $K$ quantile-based instruments $\widehat{g}_{k}(Z_{i})=Z_{i}^{\prime}\widehat{\pi}(\tau_{k})$ and let $M=[Z_{1},G]$. The projected regressor $\tilde{X}=P_{M}X$ is exactly the best linear predictor of $X$ in the span of $(Z_{1},G)$:
$\tilde{X}_{i}=Z_{1i}^{\prime}\hat{\delta}+G_{i}^{\prime}\hat{w}$, (15)
[…]
$\widehat{X}_{i}^{\mathrm{QLS}}:=\tilde{X}_{i}=P_{M}X_{i},$ (16)
then the plug-in estimator that regresses $Y$ on $(\widehat{X}^{\mathrm{QLS}},Z_{1})$ is identical (in finite samples) to the corresponding Q–LS IV estimator.
General comparison
Outside of the pure location case, Q–LS remains consistent under distributional relevance even when mean relevance fails or the conditional mean is misspecified, because it exploits shifts in the entire distribution of $X\mid Z$. In settings where both mean relevance and distributional relevance hold and the efficient IV $h^{*}(Z)$ lies in the closure of the quantile-generated class $\mathcal{H}$, Q–LS again attains the same efficiency bound as optimal 2SLS. In more general cases where $h^{*}(Z)\notin\mathcal{H}$, Q–LS is efficient within $\mathcal{H}$ but may be slightly less efficient than an oracle 2SLS that can use arbitrary functions of $Z$ as instruments. The main gain of Q–LS is robustness: it can remain strongly identified in settings where the classical first-stage mean shift is close to zero but the distribution of $X\mid Z$ varies substantially with $Z$.
5 Implementation and regularization
The population optimal Q–LS instrument $h_{\mathrm{opt}}$ solves the projection problem in Definition 4. In the finite-grid implementation, this leads to the least-squares problem (9) for the discrete weights $w\in\mathbb{R}^{K}$. When $K$ is moderate and the quantile-based instruments are highly collinear, the Gram matrix $\widehat{G}^{\prime}\widehat{G}$ can be ill-conditioned, and the unregularized solution $\widehat{w}_{n}$ becomes unstable. This is the familiar ill-posedness of inverse problems and many-instrument IV.
We discuss two regularized versions of Q–LS that address this issue: ridge-regularized Q–LS and sparse (LASSO) Q–LS.
5.1 Ridge-regularized Q–LS
A simple and effective approach is Tikhonov (ridge) regularization. Instead of (9), we solve
$\widehat{w}_{n}^{\lambda}\in\arg\min_{w\in\mathbb{R}^{K}}\left\{\frac{1}{n}\sum_{i=1}^{n}\big(X_{i}-\widehat{G}_{i}^{\prime}w\big)^{2}+\lambda_{n}\|w\|^{2}\right\},$ (17)
which yields the closed-form solution
$\widehat{w}_{n}^{\lambda}=\big(\widehat{G}^{\prime}\widehat{G}+n\lambda_{n}I_{K}\big)^{-1}\widehat{G}^{\prime}X.$ (18)
The corresponding regularized Q–LS instrument is
$\widehat{X}_{i}^{\mathrm{Q\text{--}LS},\lambda}:=\widehat{G}_{i}^{\prime}\widehat{w}_{n}^{\lambda},\qquad i=1,\dots,n.$ (19)
Finite-sample stability is improved because the ridge term $n\lambda_{n}I_{K}$ stabilizes the inversion of $\widehat{G}^{\prime}\widehat{G}$ and shrinks the weights toward zero. Under the same conditions as in Proposition 1, and provided $\lambda_{n}\to 0$ at a suitable rate (for example, $\lambda_{n}\downarrow 0$ and $K_{n}^{2}/(n\lambda_{n})\to 0$), the regularized weights $\widehat{w}_{n}^{\lambda}$ remain consistent for the population optimal weights and $\widehat{X}_{i}^{\mathrm{Q\text{--}LS},\lambda}$ converges in $L^{2}$ to $h_{\mathrm{opt}}(Z_{i})$. The asymptotic distribution in Proposition 2 therefore continues to hold with $\widehat{X}_{i}^{\mathrm{Q\text{--}LS},\lambda}$ replacing $\widehat{X}_{i}^{\mathrm{Q\text{--}LS}}$.
Two additional practical devices can be used in tandem with the ridge penalty. First, one may limit the number of quantile instruments $K$ to a moderate value (e.g. $K=5$ or $K=10$), or work with a low-dimensional principal-component representation of the columns of $\widehat{G}$. Second, instead of solving (17) explicitly, one can treat each $\widehat{g}_{k}(Z_{i})$ as a separate instrument, run 2SLS or LIML with the full instrument vector, and let the IV estimator choose an implicit linear combination of quantile-based instruments. These strategies reduce the effective instrument dimension and mitigate many-instrument bias while preserving the distributional information exploited by Q–LS.
5.2 Sparse Q–LS via LASSO
The ill-posedness of the optimal weight problem arises because the finite-grid instrument dictionary $\widehat{G}$ can be high-dimensional and highly collinear. An alternative to ridge regularization is to enforce sparsity in the weights and let the data select a subset of quantile-based instruments. This leads naturally to an $\ell_{1}$-penalized (LASSO) version of Q–LS.
Given the $K$ quantile-based instruments $\widehat{g}_{k}(Z_{i})=Z_{i}^{\prime}\widehat{\pi}_{n}(\tau_{k})$, stacked in $\widehat{G}_{i}$ and $\widehat{G}$ as in (8), we define the LASSO-regularized weights as
$\widehat{w}_{n}^{\mathrm{LASSO}}\in\arg\min_{w\in\mathbb{R}^{K}}\left\{\frac{1}{n}\sum_{i=1}^{n}\big(X_{i}-\widehat{G}_{i}^{\prime}w\big)^{2}+\lambda_{n}\|w\|_{1}\right\},$ (20)
where $\lambda_{n}>0$ is a tuning parameter. The corresponding sparse Q–LS instrument is
$\widehat{X}_{i}^{\mathrm{Q\text{--}LS,L}}:=\widehat{G}_{i}^{\prime}\widehat{w}_{n}^{\mathrm{LASSO}}=\sum_{k=1}^{K}\widehat{w}_{n,k}^{\mathrm{LASSO}}\,\widehat{g}_{k}(Z_{i}),\qquad i=1,\dots,n.$ (21)
By construction, only a subset of the quantile indexes has $\widehat{w}_{n,k}^{\mathrm{LASSO}}\neq 0$, so the effective instrument dimension is reduced and the many-instrument problem is mitigated. In practice, we also consider a post-LASSO refinement: letting $\widehat{S}_{n}=\{k:\widehat{w}_{n,k}^{\mathrm{LASSO}}\neq 0\}$ denote the selected support, we re-estimate the weights by least squares restricted to $\widehat{S}_{n}$ and use the resulting fitted values as $\widehat{X}_{i}^{\mathrm{Q\text{--}LS,L}}$.
The second-stage estimator is then defined as in (12), with $\widehat{X}^{\mathrm{Q\text{--}LS}}$ replaced by $\widehat{X}^{\mathrm{Q\text{--}LS,L}}$. From the perspective of the structural parameter $\theta_{0}$, the first stage (quantile processes and weights) is a high-dimensional nuisance. The IV moment condition is Neyman-orthogonal with respect to this nuisance, so that, under standard sparsity and rate conditions for LASSO and a suitable choice of $\lambda_{n}$, the estimation error in $\widehat{X}_{i}^{\mathrm{Q\text{--}LS,L}}$ is $o_{p}(n^{-1/2})$ in the moment equations. As a result, the asymptotic normality result in Proposition 2 continues to hold with $\widehat{X}_{i}^{\mathrm{Q\text{--}LS,L}}$ in place of $\widehat{X}_{i}^{\mathrm{Q\text{--}LS}}$, and the usual heteroskedasticity-robust 2SLS covariance matrix computed with the sparse Q–LS instrument remains valid for inference.
In empirical work, the tuning parameter $\lambda_{n}$ can be chosen either by cross-validation (targeting prediction of $X$) or by more conservative theoretical rules designed for high-dimensional linear regression. A moderate level of penalization, combined with post-LASSO refitting, tends to yield stable first stages and good finite-sample size for Q–LS inference.
5.3 Practical recommendations
We view Q–LS as a complement to, rather than a replacement for, the standard 2SLS toolkit with weak instruments. In practice, applied work can proceed in three steps: (i) run conventional mean-based diagnostics, (ii) assess distributional relevance, and (iii) choose between 2SLS and Q–LS (and its regularised variants) based on a small menu of test statistics and simple stability checks.
Step 1: Start from the usual 2SLS diagnostics.
  • Estimate the conventional first stage $X$ (endogenous variable) on $Z$ (instrument) and controls and report the usual weak-IV statistics (e.g. the cluster-robust Kleibergen–Paap $F$-statistic or the Stock–Yogo $F$ for the mean first stage).
  • If the mean first stage is clearly strong (e.g. robust $F\gtrsim 10$), treat linear 2SLS with the conventional instrument as your baseline and view Q–LS mainly as a robustness check.
  • If the mean first stage is weak (e.g. robust $F\ll 10$), proceed to distributional diagnostics rather than discarding the design.
Step 2: Diagnose distributional relevance.
  • Run a small set of conditional quantile regressions of $X$ on $Z$ and controls (e.g. $\tau\in\{0.10,0.25,0.50,0.75,0.90\}$) and test the joint significance of $Z$ in each quantile regression. Large $F$-statistics at upper quantiles with small or insignificant mean effects are indicative of purely distributional relevance.
  • Graphically, compare the estimated conditional quantiles or empirical CDFs of $X$ across instrument values (e.g. pre/post policy). Pronounced tail shifts with little change in the median or mean point towards a distributional design where Q–LS is useful.
Step 3: Choosing $K$, regularisation, and inference.
  • (i)
    Quantile grid. Use a moderate grid of conditional quantiles, for example $K\in\{5,10,20\}$ equally spaced over $(0,1)$, possibly excluding extreme tails (e.g. $\tau_{\min},1-\tau_{\min}]$ with $\tau_{\min}\in[0.02,0.05]$). In most empirical samples $K=10$ provides a good compromise between flexibility and stability.
  • (ii)
    Regularisation. If $K$ is large relative to $n$ or the quantile regressions are noisy, use ridge- or LASSO-regularised Q–LS (Q–LS–R, Q–LS–L1, Q–LS–L2) rather than the unpenalised version. Our simulations suggest that ridge (Q–LS–R) is a conservative default, while LASSO variants are helpful in very many-instrument designs.
  • (iii)
    Inference. Always report heteroskedasticity-robust (and, in panels, cluster-robust) standard errors based on the Q–LS sandwich variance (or its penalised analogue). Since Q–LS produces a conventional instrument $\hat{h}(Z)$, practitioners can also apply standard weak-IV-robust tools (e.g. Anderson–Rubin or conditional likelihood ratio tests) using $\hat{h}(Z)$ as the excluded instrument.
Step 4: Stability checks and reporting.
  • (a)
    Compare Q–LS and 2SLS. Report side-by-side Q–LS and classical 2SLS estimates and standard errors. In designs with strong mean relevance, point estimates and standard errors should be similar. In designs with weak mean relevance but strong distributional relevance, Q–LS is expected to deliver comparable point estimates with much tighter confidence intervals.
  • (b)
    Check robustness to $(K,\lambda)$. Re-estimate Q–LS on a few nearby quantile grids (e.g. $K=5,10,20$) and, for regularised variants, a small set of tuning parameters $\lambda$. Large swings in the coefficient across neighbouring grids or penalties are a warning sign that the design may be too fragile for sharp causal interpretation.
  • (c)
    Document instrument strength. Alongside the conventional first-stage $F$, report a “distributional FF”: the joint significance of the quantile-generated instruments in the Q–LS first stage. This makes transparent what Q–LS is exploiting that 2SLS misses.
Table 1 summarises these recommendations in a simple diagnosis chart.
Table 1: Rule-of-thumb diagnostics and suggested estimator
Diagnostic pattern Suggested practice
Strong mean first stage ($F_{\text{mean}}\gtrsim 10$); quantile regressions show similar shifts across quantiles Use conventional 2SLS with the mean-based instrument as the primary estimator; report Q–LS as a robustness check (expect similar point estimates and standard errors).
Weak mean first stage ($F_{\text{mean}}\ll 10$); quantile regressions show strong tail shifts (large $F$ at high quantiles) Treat the design as a distributional IV setting. Report Q–LS (and a regularised variant if $K$ is large) as the main estimator, with weak-IV-robust inference based on the Q–LS instrument; include 2SLS for comparison but emphasise its weak-IV limitations.
Both mean and distributional diagnostics weak (no clear shifts in mean or quantiles of $X$ given $Z$) Interpret any IV estimates with caution. Neither 2SLS nor Q–LS has a strong effective first stage; focus on reduced-form estimates or seek alternative instruments or sources of identification.
Q–LS and 2SLS estimates differ markedly in sign or magnitude, or Q–LS is highly sensitive to $(K,\lambda)$ Treat results as exploratory. Investigate model misspecification (functional form, outliers, covariates), reconsider the quantile specification, and report the instability explicitly rather than relying on a single specification.
6 Simulation study
This section will examine the finite-sample performance of Q–LS and its regularised variants in a set of Monte Carlo designs. We focus on three types of designs that map directly to our theoretical discussion. (A) strong mean relevance (benchmark Gaussian case), (B) Trigonometric and distributional instruments (variance shifts) and (C) mixed mean and distributional relevance.
The following parts of the data generating processes (DGPs) are the same in all three designs;
  • The structural equation
    $Y=\beta_{0}+Z_{1}\beta_{1}+X\beta_{x}+\varepsilon$
    where the parameter on the endogenous variable $X$, $\beta_{x}=1$ is the parameter of interest, and $\beta_{0}=1,\beta_{1}=1$ with $Z_{1}\sim N(0,1)$.
  • The errors between the structural equation ($\varepsilon$) and first-stage ($\nu$) follow the following process,
    $(\varepsilon,\nu)$ $\sim N\left(0,\begin{pmatrix}1&0.6\\ 0.6&1\end{pmatrix}\right)$.
  • We will consider sample sizes representative of typical applied work $n=500,1000$ and we run 1000 replications.
We compare 10 estimators:
  1. 1.
    Classical IV where the first stage is a least squares projection with just first order terms i.e.
    $X$ $=\zeta_{0}+\zeta_{1}Z_{1}+\zeta_{2}Z_{2}+e_{x}$ (22)
  2. 2.
    Our Q–LS with all quantiles equally weighted, first stage is estimated as a second order polynomial,
    $X$ $=\zeta_{\tau_{k,0}}+\zeta_{\tau_{k,1}}Z_{1}+\zeta_{\tau_{k,2}}Z_{2}+\zeta_{\tau_{k,3}}Z_{1}^{2}+\zeta_{\tau_{k,4}}Z_{2}^{2}+\zeta_{\tau_{k,5}}Z_{1}Z_{2}+e_{x,\tau_{k}}$ (23)
    for a grid of quantiles $\{\tau_{k}\}$. The resulting fitted quantile regression functions are used to generate a prediction of the endogenous variable by an equally weighted average of the quantile regression fitted values,
    1. a.
      Estimate of $\beta_{x}$ is from analytical 2SLS analogue formulas (12) denoted Q–LS-a.
    2. b.
      Estimate of $\beta_{x}$ is from a plug-in OLS regression of the structural equation where the equally weighted average of quantile fitted values is used in place of $X$, denoted Q–LS.
  3. 3.
    Distributionally relevant control function, first stage is also estimated as a second order polynomial (23) and the control function is estimates as in Chernozhukov et al. (2020).
    $y$ $=\beta_{0}+\beta_{1}z_{1}+\beta_{x}x+\beta_{v}\hat{v}+e_{y}$ (24)
    where $\hat{v}\approx F_{X}(X|Z)$ is a control function and is defined in (35). Estimate of $\beta_{x}$ is from (24), denoted DR-IV.
  4. 4.
    Our Q–LS with all quantiles weighted by ridge regression coefficients, Same as estimator (2) but where the weights are ridge regression coefficients.
    1. a.
      Estimate of $\beta_{x}$ is from analytical 2–SLS analogue formulas (12) denoted Q–LS-R-a.
    2. b.
      Estimate of $\beta_{x}$ is from a plug-in OLS regression of the structural equation, where the ridge weighted average of quantile fitted values is used in place of $X$, denoted Q–LS-R.
  5. 5.
    Our Q–LS with all LASSO selected quantiles are weighted equally weighted (all other given zero weights), Same as estimator (2) but only LASSO selected quantile fitted values given non-zero weights.
    1. a.
      Estimate of $\beta_{x}$ is from analytical 2–SLS analogue formulas (12) denoted Q–LS-L1-a.
    2. b.
      Estimate of $\beta_{x}$ is from a plug-in OLS regression of the structural equation where the LASSO selected weighted average of quantile fitted values is used in place of $X$ denoted Q–LS-L1.
  6. 6.
    Our Q–LS with all quantiles weighted by LASSO regression coefficients, Same as estimator (2) but where the weights are LASSO regression coefficients.
    1. a.
      Estimate of $\beta_{x}$ is from analytical 2–SLS formulas in the main paper denoted Q–LS-L2-a.
    2. b.
      Estimate of $\beta_{x}$ is from a plug-in OLS regression of the structural equation where the LASSO weighted average of quantile fitted values is used in place of $X$ denoted Q–LS-L2.
The weights for the weighted average of the quantiles are generated by sample splitting. The first 50% of the data is used to estimate the Ridge/LASSO regression regularisation parameters with 10 fold cross-validation predication accuracy. The estimated hyperparamter and the other 50% of the data is then used to estimated the Ridge/LASSO coefficient estimate.
For all Q–LS variants and DR-CF we consider quantiles between $v=0.01$ and 1v1-v and set the gap between the quantiles to $c=\{0.10,0.05,0.01,0.001\}$ this corresponds to using $K=\{10,20,99,981\}$ quantiles to generate the average of the quantile regression fitted values of $X$ or approximate the CDF ($\hat{v}$). The trimming of the quantiles follows Chernozhukov et al. (2020).
Simulation results will be reported in terms of bias, root mean squared error (RMSE), and empirical 95% coverage of the simulation estimate.
Design A: Strong mean relevance (benchmark Gaussian case)
The first design is a “best-case” benchmark in which classical mean relevance holds and the conditional distribution of $X\mid Z$ belongs to a Gaussian location family. The data-generating process of the first stage is
$X$ $=\alpha_{1}Z_{1}+\alpha_{2}Z_{2}+\nu$
$Z_{2}$ $\sim\mathcal{N}(0,1)$
with $\alpha_{1}=\alpha_{2}=1$.
In this design, both 2SLS and Q–LS are correctly specified and the conditions of Proposition 3 hold. The simulation will document:
  • similarity of bias, RMSE and empirical coverage for Q–LS and 2–SLS;
  • sensitivity of Q–LS to the choice of quantile grid and regularisation in a setting where both estimators are efficient.
Design B: Trigonometric and distributional instruments
The second design studies a case where the excluded instrument $Z_{2}$ generates rich non-linear and distributional variation in $X$, while the linear first stage of $X$ on $Z_{2}$ can be arbitrarily weak. The data-generating process of the first stage is
$X$ $=\alpha_{1}Z_{1}-\cos(\omega Z_{2})+\exp(\gamma Z_{2})\nu$ (25)
$Z_{2}$ $\sim\mathcal{N}(0,1)$
with $\alpha_{1}=1$ and a variance parameter $\gamma=0.5$.
Hence, conditional on $(Z_{1},Z_{2})$,
$\mathbb{E}[X\mid Z_{1},Z_{2}]=\alpha_{1}Z_{1}-\cos(\omega Z_{2}),\qquad\operatorname{Var}(X\mid Z_{1},Z_{2})=\exp(2\gamma Z_{2}).$
The excluded instrument $Z_{2}$ affects the conditional mean through the nonlinear term $\cos(\omega Z_{2})$ and the conditional dispersion through $\exp(\gamma Z_{2})$. The conditional distribution $F_{X\mid Z}(\cdot\mid Z_{1},Z_{2})$ is therefore strongly distributionally relevant in the sense of the main text.
Note that, since $Z_{2}\sim\mathcal{N}(0,1)$ is symmetric around zero and $z\mapsto\cos(\omega z)$ is an even function, we have
$Cov\big(Z_{2},\cos(\omega Z_{2})\big)=\mathbb{E}\big[-Z_{2}\cos(\omega Z_{2})\big]=0$
for all $\omega\geq 0$. Moreover, $\nu$ is independent of $Z_{2}$ and has mean zero, so $Cov(Z_{2},\exp(\gamma Z_{2})\nu)=0$. Thus
$Cov(X,Z_{2})=0$
in the population for every value of $\omega$. Classical linear IV based on the regression of $X$ on $Z_{2}$ therefore has an exactly zero linear first stage, even though $Z_{2}$ induces strong nonlinear and distributional variation in $X$.
Sub-designs B1–B3.
We consider three values of the frequency parameter $\omega$:
  • Design B1 ($\omega=0$).
    $X=\alpha_{1}Z_{1}+\exp(\gamma Z_{2})\,\nu.$
    The conditional mean is linear in $Z_{1}$ and does not depend on $Z_{2}$,
    $\mathbb{E}[X\mid Z_{1},Z_{2}]=\alpha_{1}Z_{1},$
    so the excluded instrument $Z_{2}$ is purely distributional: it affects only the variance of $X$ through $\exp(\gamma Z_{2})$, and $\mathbb{E}[X\mid Z_{2}]$ is constant. Classical 2SLS that treats $Z_{2}$ as a mean shifter is therefore weak, while Q–LS can exploit the variance shifts.
  • Design B2 ($\omega=0.01$).
    $X=\alpha_{1}Z_{1}-\cos(0.01Z_{2})+\exp(\gamma Z_{2})\,\nu.$
    Here
    $\mathbb{E}[X\mid Z_{2}]=-\cos(0.01Z_{2})$
    is non-constant but very flat. The linear projection of $X$ on $Z_{2}$ has a small slope and a low first-stage $F$-statistic, so classical 2SLS using $Z_{2}$ appears weak. Nevertheless, the true mean and the conditional distribution depend nontrivially on $Z_{2}$, so the optimal Q–LS instrument $h_{\mathrm{opt}}(Z)$ is non-degenerate and remains strongly relevant.
  • Design B3 ($\omega=0.5$).
    $X=\alpha_{1}Z_{1}-\cos(0.5Z_{2})+\exp(\gamma Z_{2})\,\nu.$
    The conditional mean
    $\mathbb{E}[X\mid Z_{2}]=-\cos(0.5Z_{2})$
    is highly nonlinear and oscillatory. As $\omega$ increases, the covariance $\mathrm{Cov}(X,Z_{2})$, and hence the linear first-stage coefficient of $X$ on $Z_{2}$, becomes small even though the amplitude of $\cos(\omega Z_{2})$ remains large. Thus the linear first stage for 2SLS can look weak or misleading, while the quantile-generated space used by Q–LS can approximate $\cos(0.5Z_{2})$ well, delivering a strong and stable first stage in the Q–LS sense.
Overall, Designs B1–B3 trace a progression from a purely distributional instrument (B1, no mean effect of $Z_{2}$) through an “almost mean-irrelevant” case (B2, very flat mean effect) to a design with strong nonlinear mean and variance effects (B3). This family highlights that classical first-stage diagnostics based on the linear regression of $X$ on $Z_{2}$ can substantially understate instrument strength, whereas Q–LS remains well identified whenever $X$ has a nonzero projection in the quantile-generated space.
Design C: Mixed mean and distributional relevance
The third design introduces both mean and variance effects:
  • $Z_{2}\in\{0,1\}$ with $\mathbb{P}(Z_{2}=0)=\mathbb{P}(Z_{2}=1)=1/2$.
  • $X=\alpha_{1}Z_{1}+\alpha_{2}Z_{2}+\sigma(Z_{2})\nu$, with $\sigma(0)=1$, $\sigma(1)=3$, so that $\mathbb{E}[X\mid Z_{2}]\neq 0$, and $\alpha_{1}=\alpha_{2}=1$.
Here, both mean relevance and distributional relevance hold, but the contribution of each can be tuned by varying $(\alpha_{2},\sigma(0),\sigma(1))$. The simulation will:
  • study the performance of Q–LS and 2SLS as the mean effect $\alpha_{2}$ becomes small while variance shifts remain large;
  • assess whether Q–LS remains competitive with 2SLS when mean effects are strong, and dominates when mean effects are weak but distributional shifts are large.
Note as $Z_{2}$ is binary in this design we set $\zeta_{\tau_{k,4}}=0$ in (23) to avoid a singular QR design matrix, for all Q–LS variants and DR-CF.
Simulation Results
Table 2: Design A and C finite-sample performance
Design A Design C
K Estimator Bias(RMSE)Cov. Bias(RMSE)Cov. Bias(RMSE)Cov. Bias(RMSE)Cov.
Classic IV 0.001(0.046)0.942 -0.001(0.032)0.946 -0.010(0.094)0.961 -0.012(0.070)0.957
10 DR-CF 0.046(0.063)0.689 0.038(0.048)0.623 0.058(0.066)0.498 0.053(0.058)0.312
10 Q–LS 0.003(0.048)0.999 0.001(0.033)1.000 0.615(0.686)0.878 0.663(0.718)0.377
10 Q–LS-a 0.004(0.046)0.937 0.001(0.031)0.946 0.021(0.095)0.928 -0.560(17.875)0.934
10 Q–LS-R 0.005(0.049)0.999 0.002(0.033)0.999 0.042(0.272)0.966 0.023(0.189)0.962
10 Q–LS-R-a 0.004(0.046)0.939 0.001(0.031)0.946 0.010(0.087)0.941 -0.001(0.068)0.942
10 Q–LS-L1 0.002(0.057)0.992 0.000(0.040)0.994 0.018(0.271)0.957 0.011(0.193)0.952
10 Q–LS-L1-a 0.004(0.046)0.938 0.001(0.031)0.946 0.010(0.087)0.942 -0.001(0.068)0.941
20 DR-CF 0.026(0.051)0.818 0.018(0.035)0.828 0.032(0.047)0.773 0.027(0.036)0.742
20 Q–LS 0.004(0.047)1.000 0.001(0.032)0.999 0.192(0.220)1.000 0.187(0.202)1.000
20 Q–LS-a 0.004(0.046)0.938 0.001(0.031)0.946 0.012(0.088)0.942 0.000(0.067)0.939
20 Q–LS-R 0.005(0.047)1.000 0.001(0.032)0.999 0.036(0.264)0.967 0.018(0.186)0.966
20 Q–LS-R-a 0.004(0.046)0.938 0.001(0.031)0.946 0.010(0.087)0.941 -0.001(0.068)0.942
20 Q–LS-L1 0.002(0.057)0.991 0.000(0.041)0.989 0.019(0.269)0.958 0.011(0.192)0.958
20 Q–LS-L1-a 0.004(0.046)0.939 0.001(0.031)0.946 0.010(0.087)0.940 -0.001(0.068)0.942
99 DR-CF 0.015(0.048)0.849 0.008(0.032)0.858 0.018(0.040)0.873 0.011(0.028)0.880
99 Q–LS 0.004(0.046)1.000 0.001(0.032)0.999 0.015(0.092)1.000 0.001(0.069)1.000
99 Q–LS-a 0.004(0.046)0.938 0.001(0.031)0.946 0.009(0.088)0.943 -0.002(0.067)0.945
99 Q–LS-R 0.006(0.048)0.999 0.001(0.033)0.998 0.030(0.266)0.966 0.014(0.189)0.965
99 Q–LS-R-a 0.004(0.046)0.937 0.001(0.031)0.945 0.010(0.088)0.941 -0.001(0.068)0.943
99 Q–LS-L1 0.004(0.060)0.991 0.002(0.045)0.982 0.015(0.264)0.959 0.011(0.192)0.958
99 Q–LS-L1-a 0.004(0.046)0.940 0.001(0.032)0.946 0.011(0.087)0.940 -0.001(0.068)0.941
981 DR-CF 0.015(0.047)0.849 0.007(0.032)0.858 0.016(0.039)0.880 0.010(0.028)0.885
981 Q–LS 0.004(0.046)1.000 0.001(0.032)0.999 0.015(0.092)1.000 0.001(0.069)1.000
981 Q–LS-a 0.004(0.046)0.938 0.001(0.031)0.946 0.009(0.088)0.944 -0.002(0.067)0.945
981 Q–LS-R 0.004(0.052)0.997 0.000(0.037)0.998 0.001(0.252)0.959 -0.009(0.184)0.956
981 Q–LS-R-a 0.004(0.046)0.939 0.001(0.032)0.946 0.009(0.087)0.943 -0.002(0.068)0.944
981 Q–LS-L1 0.003(0.062)0.989 0.002(0.046)0.977 0.016(0.264)0.959 0.012(0.192)0.959
981 Q–LS-L1-a 0.004(0.046)0.940 0.001(0.031)0.946 0.011(0.087)0.942 -0.001(0.068)0.943
n=500 n=1000 n=500 n=1000
Note: Bias and RMSE are Monte Carlo averages over 1,000 replications. “Cov.” is the empirical coverage of nominal 95% confidence intervals. K is the number of quantiles used. The full results, including average asymptotic standard error and simulation standard deviation, and Q–LS-L2 are available upon request.
Table 3: Design B finite-sample performance
K Estimator Bias(RMSE)Cov. Bias(RMSE)Cov. Bias(RMSE)Cov. Bias(RMSE)Cov.
Classic IV 0.199(5.745)0.989 0.627(25.900)0.988 -1.063(34.095)0.993 2.187(67.609)0.987
10 DR-CF 0.122(0.135)0.386 0.106(0.113)0.209 0.116(0.129)0.412 0.100(0.107)0.234
10 Q–LS -0.127(0.937)0.991 -0.638(0.979)0.961 0.007(0.446)0.982 -0.167(0.390)0.930
10 Q–LS-a 0.026(6.982)0.954 1.297(27.315)0.977 0.121(0.627)0.904 0.064(0.291)0.932
10 Q–LS-R 0.212(0.910)0.982 0.162(1.087)0.987 0.003(0.437)0.989 -0.046(0.361)0.982
10 Q–LS-R-a 0.318(6.496)0.937 5.713(156.401)0.952 0.056(1.340)0.899 -0.031(2.164)0.934
10 Q–LS-L1 -0.008(0.794)0.984 -0.027(0.961)0.980 -0.067(0.443)0.986 -0.083(0.375)0.979
10 Q–LS-L1-a 1.366(29.108)0.932 0.730(18.423)0.959 0.063(1.258)0.907 0.068(0.995)0.938
20 DR-CF 0.075(0.097)0.717 0.057(0.070)0.694 0.071(0.094)0.716 0.053(0.067)0.708
20 Q–LS 0.544(0.989)0.992 0.268(0.953)0.988 0.184(0.445)0.992 0.079(0.335)0.991
20 Q–LS-a 0.539(4.909)0.915 0.472(7.595)0.951 0.113(0.564)0.884 0.052(0.227)0.934
20 Q–LS-R 0.222(0.915)0.984 0.179(0.960)0.988 0.000(0.441)0.987 -0.054(0.359)0.983
20 Q–LS-R-a -0.037(10.468)0.936 0.185(4.615)0.954 0.094(0.576)0.903 0.136(3.232)0.932
20 Q–LS-L1 -0.023(0.783)0.979 -0.048(0.919)0.983 -0.068(0.445)0.985 -0.083(0.369)0.982
20 Q–LS-L1-a 0.386(5.726)0.934 3.381(93.580)0.958 0.088(1.120)0.907 0.034(0.365)0.937
99 DR-CF 0.045(0.078)0.835 0.027(0.052)0.861 0.043(0.077)0.830 0.025(0.049)0.867
99 Q–LS 0.744(1.045)0.990 0.714(1.068)0.993 0.170(0.438)0.996 0.057(0.325)0.993
99 Q–LS-a 0.493(4.270)0.910 0.289(3.596)0.942 0.077(0.729)0.889 0.036(0.234)0.931
99 Q–LS-R 0.205(0.929)0.982 0.134(0.953)0.988 -0.009(0.447)0.987 -0.064(0.363)0.984
99 Q–LS-R-a 2.639(54.067)0.930 -0.033(14.738)0.960 0.081(0.630)0.901 0.109(2.470)0.936
99 Q–LS-L1 -0.050(0.759)0.981 -0.061(0.894)0.983 -0.078(0.442)0.985 -0.086(0.373)0.980
99 Q–LS-L1-a 0.398(2.288)0.933 0.212(8.797)0.957 0.073(1.885)0.908 0.036(0.374)0.933
981 DR-CF 0.044(0.077)0.848 0.025(0.051)0.877 0.041(0.076)0.839 0.023(0.049)0.876
981 Q–LS 0.742(1.045)0.987 0.701(1.108)0.993 0.171(0.439)0.996 0.057(0.326)0.993
981 Q–LS-a 0.355(0.725)0.910 0.497(2.448)0.940 0.076(0.764)0.890 0.036(0.234)0.931
981 Q–LS-R 0.200(0.917)0.984 0.119(0.949)0.986 -0.023(0.440)0.988 -0.075(0.363)0.982
981 Q–LS-R-a 0.362(3.965)0.935 0.554(10.794)0.955 0.008(3.065)0.899 0.037(0.429)0.935
981 Q–LS-L1 -0.052(0.773)0.982 -0.064(0.908)0.986 -0.079(0.438)0.985 -0.085(0.369)0.979
981 Q–LS-L1-a 0.414(6.936)0.933 1.274(30.298)0.955 0.119(0.967)0.909 0.036(0.375)0.931
n=500 & $\omega$=0 n=1000 & $\omega$=0 n=500 & $\omega$=0.5 n=1000 & $\omega$=0.5
Note: Bias and RMSE are Monte Carlo averages over 1,000 replications. “Cov.” is the empirical coverage of nominal 95% confidence intervals. K is the number of quantiles used. The full results, including average asymptotic standard error and simulation standard deviation, and Q–LS-L2 are available upon request.
Table 2 presents the results for Design A and Design C for sample sizes of 500 and 1000.555Full results for Design A and Design C are available upon request. It is important to note that classical IV is correctly specified, as $Z_{2}$ is strongly mean-relevant in both Design A and Design C. For Design A, we find that all Q-LS variants perform similarly to classical IV, regardless of the number of quantiles used. This is due to the quantile projection being similar to the mean projection. The analytical Q-LS variants provide better coverage (closer to 0.95) than the OLS plug-in standard errors, which tend to yield over-coverage, indicating conservative standard errors in this design. The control function version (DR-CF) performs worse in terms of bias, RMSE, and coverage than all Q-LS variants and classical IV, regardless of sample size. However, its performance does improve as the number of quantiles increases in this design.
For Design C, $Z_{2}$ is both strongly mean-relevant and variance-relevant. As mean relevance is less clean, the bias and RMSE of classical IV are ten and two times as large, respectively, as in Design A for a given sample size. Unlike Design A, the unregularized Q-LS variants exhibit substantial bias when 10 quantiles are used; however, this bias decreases as the number of quantiles and the sample size increase. When 981 quantiles are used, all Q-LS variants have the same or smaller bias than classical IV, and the regularized versions (plug-in and analytical) all yield good coverage. The performance of DR-CF relative to the other estimators and as the number of quantiles and the sample size increase, is similar to that in Design A.
Table 3 shows the results for Design B with $\omega={0,0.5}$. [^6] Full results for Design B are available upon request. Interestingly, classical IV performs worse when $\omega$ is non-zero in terms of both bias and RMSE for a given sample size, whereas the bias and RMSE of the Q-LS variants decrease, except for the bias of Q–LS-L1, which marginally increases (in absolute value) as the number of quantiles increases. More quantiles are needed for better approximation because the distributional relevance of $Z_{2}$ becomes stronger as $\omega$ increases due to the $\cos$ transformation. Similar to Designs A and C, DR-CF performance improves as the number of quantiles increases. However, unlike Designs A and C, DR-CF has the smallest bias and RMSE when $\omega=0$ for a given sample size, and also the smallest bias and RMSE when $\omega=0.5$ and $n=1000$. Despite this, DR-CF coverage is the worst among all estimators considered for a given number of quantiles and sample size.
Overall, across all designs considered, we find that at least one variant of Q-LS with 99 or more quantiles performs as well as or better than classical IV, even in designs where the instrument $Z_{2}$ is strongly mean-relevant. When a large number of quantiles are used, analytical Q-LS empirical coverage tends to be close to, but slightly below, 0.95, while the plug-in variants tend to exhibit coverage above 0.95. When the instrument is not strongly mean-relevant, the plug-in point estimates tend to be more stable than their analytical counterparts. The DR-CF estimator performs better as the number of quantiles increases across all designs considered, but it requires a large number of quantiles and weak mean relevance to achieve lower bias and RMSE than all Q-LS variants.
In Design A and C (strong mean relevance), Q–LS essentially reproduces optimal 2SLS. This confirms the theoretical prediction that, in a pure location-shift setting with strong instruments, exploiting the full conditional quantile process does not alter first-order behavior. Design B1 illustrates the degenerate branch of Lemma 1. Because the optimal Q–LS instrument $h_{\mathrm{opt}}(Z)$ is identically zero, the first stage for Q–LS collapses and the estimator fails in the usual weak-IV sense. We do not regard this as a pathology of Q–LS, but as a reminder that distributional relevance alone is not sufficient for a useful first stage; Assumption 1 is needed to rule out such cases. The main action comes in Design B3, where instruments are weak in the mean but strong in the distribution. Here classical 2SLS exhibits substantial finite-sample distortions. In contrast, Q–LS estimators based on a moderate grid of quantiles (e.g. $K=10$) are well centered and display near-correct coverage across the same sample sizes. These patterns mirror the HRS application: when policy reforms primarily reshape tails and dispersion rather than means, exploiting distributional relevance through Q–LS restores a strong, stable first stage from the same underlying instrument.
7 Empirical application
We use the Health and Retirement Study (HRS) to study how financial exposure to out-of-pocket (OOP) medical spending risk affects mental health among older Americans. Our main outcome is an indicator for any depressive symptoms (CESD) among respondents aged 65 and older. The endogenous regressor is a scalar measure of annual OOP medical spending, and the key policy instrument is an indicator for the post–Medicare Part D period (year $\geq 2006$), during which subsidized prescription drug coverage expanded and catastrophic protection was introduced.
To mirror the simulation designs, we work with two complementary specifications. First, we treat OOP spending in nominal dollars and use a single post–Part D indicator as the excluded instrument. This “stress test” design generates a classical weak-IV environment when the risk index is summarized by mean OOP spending, and it provides a direct counterpart to the distributional simulation in Section 6. Second, we express OOP spending in real 2015 dollars using the OECD CPI and show that the same policy indicator becomes a strong instrument in this scaling. The real-dollar specification serves as a well-identified benchmark in which Q–LS and conventional 2SLS can be expected to coincide when we omit a time trend. Once we add a linear trend in survey year, the point estimates remain very similar but Q–LS delivers somewhat tighter standard errors. Throughout, we trim the top 1% of the OOP distribution to reduce the influence of extreme outliers.
7.1 Descriptive evidence on depression and medical spending risk
Figure 1: Mean out-of-pocket medical spending and depression prevalence among HRS respondents aged 65+. Notes: The figure plots mean annual out-of-pocket (OOP) medical spending (thousands of dollars, left axis) and the share of respondents with any depressive symptoms (CESD indicator, right axis) by survey year for HRS respondents aged 65 and older. The vertical line at 2006 marks the introduction of Medicare Part D.
Descriptive patterns.
Figure 1 plots average out-of-pocket (OOP) medical spending in thousands of dollars (left axis) and the share of respondents with any depressive symptoms (right axis) among HRS respondents aged 65 and older from 2000 to 2014. Mean OOP spending rises from about \$3,000 in 2000 to a peak of roughly \$5,200 in the mid-2000s, then falls back toward \$4,000–\$4,300 in the late 2000s and early 2010s. Over the same period, the prevalence of depression declines fairly steadily: the share with any depression falls from around 19–20 percent in 2000 to roughly 11–12 percent by the early 2010s.
These raw trends are consistent with the idea that older households faced substantial OOP risk in the early 2000s, followed by some compression of spending risk—coinciding with the implementation of Medicare Part D in 2006—while mental health outcomes improved over time. However, the figure is purely descriptive: it does not control for changes in the composition of the 65+ population, secular improvements in treatment, or other contemporaneous policy changes. In the regression analysis below, we therefore focus on isolating the effect of changes in financial risk exposure, as measured by OOP spending, using instrumental‐variables methods. In addition, we trim the regression sample at the 99th percentile of realized annual OOP spending to remove a small number of extreme outliers; this has negligible impact on the aggregate time-series patterns in Figure 1 but stabilizes the covariate–outcome relationships.
Figure 2: Mean and upper quantiles of out-of-pocket medical spending by year (HRS 65+). Notes: The figure plots the mean, median (P50), and upper quantiles (P75, P90, P95) of annual OOP medical spending for HRS respondents aged 65 and older, from 2000 to 2020. The vertical line at 2006 marks the introduction of Medicare Part D. While the median and P75 remain relatively flat over time, the upper quantiles rise in the early 2000s and then flatten or decline after 2006, indicating a post–Part D compression of the right tail of the OOP spending distribution.
Figure 2 plots the mean, median, and upper quantiles (75th, 90th, and 95th percentiles) of annual OOP medical spending for HRS respondents aged 65 and older between 2000 and 2020. The median and 75th percentile are remarkably stable over time, indicating that typical OOP spending changes little over this period. By contrast, the mean closely tracks movements in the upper tail: the 90th and 95th percentiles rise sharply in the early 2000s and then flatten or decline following the introduction of Medicare Part D in 2006. The resulting compression of the gap between the upper quantiles and the median after 2006 is consistent with Part D primarily reducing exposure to extreme spending realizations rather than shifting the center of the distribution. From the perspective of our identification strategy, this pattern illustrates why standard first-stage diagnostics based on mean shifts label the post–Part D indicator as a weak instrument for mean OOP risk, even though it is strongly relevant for the distribution of spending, especially in the right tail.
7.2 Nominal OOP and a weak-IV design
We begin with a deliberately “stressful” specification that uses nominal OOP medical spending (in thousands of dollars) as the endogenous regressor and a single post–Part D indicator as the excluded instrument. This design closely mirrors the weak-IV, heavy-tailed environments in our Monte Carlo exercises and is useful for showing how Q–LS behaves when the instrument is weak for the mean but strongly relevant for the distribution of OOP spending.
Table 4 reports linear probability models relating an indicator for any depressive symptoms to mean OOP medical spending. The sample consists of 16,313 person–year observations for respondents aged 65 and older in the RAND HRS panel between 2000 and 2020, after excluding person–years with OOP spending above the 99th percentile of the nominal OOP distribution.
Table 4: Depression and financial exposure to medical spending (HRS 65+)
(1) (2)
OLS 2SLS
OOP medical spending (k\$) [oop_med_k][oop\_med\_k] $0.002^{***}$ $0.442$
$(0.001)$ $(0.410)$
Age $001^{**} 0.0$13-0.013
$(0.000)$ $(0.013)$
Medicaid $0.032^{***}$ 0.224-0.224
$(0.005)$ $(0.181)$
Medicare 0.156-0.156 0.943-0.943
$(0.098)$ $(0.741)$
Sex (female=1) $040^{***} 0.0$450.045
$(0.006)$ $(0.084)$
Hispanic $0.094^{***}$ $0.441$
$(0.011)$ $(0.330)$
Observations $16313$ $16313$
Clustered standard errors in parentheses.
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Notes: The dependent variable is an indicator for any depressive symptoms (CESD). The main regressor is nominal out-of-pocket medical spending in thousands of dollars (oop_med_koop\_med\_k). The sample consists of 16,313 person–year observations for HRS respondents aged 65 and older between 2000 and 2020, excluding the top 1% of the nominal OOP spending distribution. All regressions control for age, Medicaid coverage, Medicare coverage, sex, and Hispanic ethnicity (shispan), and include a constant. Column (2) reports 2SLS estimates where oop_med_koop\_med\_k is instrumented with an indicator for the post–Medicare Part D period (year $\geq 2006$). Standard errors are clustered at the individual level.
In column (1), higher nominal OOP spending is positively and statistically significantly associated with depression: a \$1,000 increase in annual OOP spending is associated with a 0.2 percentage point increase in the probability of reporting depressive symptoms ($\hat{\beta}=0.002$, s.e. 0.001). The 2SLS specification in column (2), which instruments mean OOP spending with post–Part D, yields a much larger point estimate ($\hat{\beta}=0.442$) but an even larger standard error (s.e. 0.410). The IV estimate is therefore imprecise and not statistically different from zero at conventional levels.
Table 5 shows that this lack of precision reflects the fact that the post–Part D indicator is a very weak instrument for mean nominal OOP spending in the trimmed sample: the cluster-robust partial $F$-statistic for the excluded instrument is 1.19, far below usual thresholds.
Table 5: First-stage regression and weak-IV test (mean nominal OOP spending)
OOP medical spending (k\$)
Post–Part D indicator 0.083-0.083
$(0.076)$
Age $0.032^{***}$
$(0.006)$
Medicaid $0.446^{***}$
$(0.069)$
Medicare $1.780^{***}$
$(0.215)$
Sex (female=1) $0.193^{**}$
$(0.075)$
Hispanic $0.773^{***}$
$(0.098)$
Observations $16313$
Weak-IV $F$ (Post–Part D) $1.19$
p-value $0.275$
Standard errors in parentheses.
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Notes: The dependent variable is nominal OOP medical spending in thousands of dollars (oop_med_koop\_med\_k). The regression includes age, Medicaid coverage, Medicare coverage, sex, and Hispanic ethnicity, as well as a constant. The sample consists of 16,313 person–year observations for HRS respondents aged 65 and older between 2000 and 2020, excluding the top 1% of the nominal OOP spending distribution. Standard errors are clustered at the individual level. “Weak-IV $F$ (Post–Part D)” reports the cluster-robust first-stage $F$-statistic for the excluded instrument from the corresponding 2SLS specification in Table 4, along with its $p$-value.
Q–LS vs 2SLS in the weak nominal design.
We now use this weak-IV nominal specification as a stress test to compare Q–LS and linear 2SLS. Our goal here is diagnostic: to document how the two estimators behave when the instrument is weak for the mean, yet strongly relevant for the distribution of OOP spending.
Table 6 compares the Q–LS estimate based on the projected OOP risk index with the conventional linear 2SLS estimate that uses mean nominal OOP spending as the endogenous regressor.
Table 6: Q–LS vs linear 2SLS: effect of nominal OOP risk on depression
(1) (2)
Q–LS 2SLS
Projected nominal OOP (Q–LS index, k\$) $0.442^{***}$
$(0.071)$
Nominal OOP spending (k\$) $0.442$
$(0.410)$
Observations $16313$ $16313$
Standard errors in parentheses.
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Notes: The dependent variable is an indicator for any depressive symptoms (CESD). Column (1) reports the Q–LS estimate obtained by regressing depression on the projected nominal OOP index $x_{\mathrm{QLS}}$ (in thousands of dollars) and controls, treating $x_{\mathrm{QLS}}$ as a generated regressor and using cluster-robust standard errors. Column (2) reports the conventional 2SLS estimate using nominal mean OOP spending as the endogenous regressor instrumented with the post–Part D indicator. All specifications control for age, Medicaid coverage, Medicare coverage, sex, and Hispanic ethnicity, and include a constant. The sample consists of 16,313 person–year observations between 2000 and 2020, trimming the top 1% of the nominal OOP spending distribution. Standard errors are clustered at the individual level.
The point estimates in columns (1) and (2) are numerically identical (0.442), but the standard errors differ dramatically. In the weak-IV nominal specification, Q–LS delivers a standard error of 0.071 and a precisely estimated positive effect, whereas linear 2SLS yields a standard error of 0.410 and a statistically insignificant estimate. This pattern mirrors our simulation results: when the policy instrument reshapes the distribution of OOP spending but produces only a noisy mean shift, Q–LS can extract a much stronger effective first stage from the same variation, tightening confidence intervals without changing the underlying signal.
In this nominal specification, the combination of weak mean relevance and heavy tails also makes the estimated magnitude difficult to interpret substantively. In the next subsection, we therefore turn to a more natural specification that expresses OOP spending in real terms.
7.3 Real OOP: from strong-IV benchmark to trend-adjusted design
For substantive interpretation, it is more natural to express OOP spending in real dollars. In the second part of the empirical section we therefore: (i) redefine OOP spending in thousands of 2015 U.S. dollars, and (ii) show how the Part D instrument looks first like a strong mean instrument, and then, after soaking up secular trends, like a primarily distributional instrument where Q–LS mainly improves precision.
We deflate nominal OOP using the OECD consumer price index obtained from FRED, constructing a monthly index with base year 2015 and aggregating to the relevant survey years. Real OOP spending in thousands of 2015 dollars is denoted oop_med_2015_koop\_med\_2015\_k. [^7] Organization for Economic Co-operation and Development, Consumer Price Indices (CPIs, HICPs), COICOP 1999: Consumer Price Index: Total for United States, retrieved from FRED, Federal Reserve Bank of St. Louis; https://fred.stlouisfed.org/series/USACPIALLMINMEI, accessed January 14, 2026.
Real OOP without time trend: a strong-IV benchmark.
In the first real-dollar specification, we repeat the nominal regressions replacing oop_med_koop\_med\_k with oop_med_2015_koop\_med\_2015\_k and keeping the same controls. In this case, the post–Part D indicator is a strong instrument for real OOP: the first-stage partial $F$-statistic is around 50 in the trimmed sample. Table 7 compares Q–LS and linear 2SLS for the depression outcome in this strong-IV environment.
Table 7: Q–LS vs linear 2SLS: effect of real OOP risk on depression (2015\$)
(1) (2)
Q–LS 2SLS
Projected real OOP (Q–LS index, 2015 k\$) $0.045^{***}$
$(0.007)$
Real OOP spending (2015 k\$) $0.045^{***}$
$(0.008)$
Observations $16311$ $16311$
Standard errors in parentheses.
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Notes: The dependent variable is an indicator for any depressive symptoms (CESD). In column (1), the endogenous regressor is the projected real OOP index in thousands of 2015 dollars constructed via the Q–LS procedure; the coefficient is estimated by OLS with cluster-robust standard errors. In column (2), real OOP spending in thousands of 2015 dollars is instrumented with the post–Part D indicator, and the coefficient is estimated by linear 2SLS with cluster-robust standard errors. All regressions control for age, Medicaid, Medicare, gender, and Hispanic origin. OOP spending is deflated to 2015 dollars using the OECD CPI (index 2015=100), and the sample trims the top 1% of the real OOP distribution. Standard errors are clustered at the individual level.
In this specification, Q–LS and 2SLS are virtually indistinguishable: both yield an effect of about 0.045 per \$1,000 of real OOP, with very similar standard errors. This is exactly the “well behaved” case in which our theory predicts that Q–LS collapses to optimal 2SLS when the instrument is strong and the linear mean representation is adequate. Table 10 , in Appendix, confirms that three different ways of evaluating uncertainty for the Q–LS coefficient—cluster-robust OLS, clustered bootstrap, and the analytical Q–LS sandwich formula—deliver nearly identical standard errors.
Real OOP with time trend: a distributional relevance design.
Finally, we add a flexible linear time trend to the real-OOP specification. This absorbs much of the secular drift in spending and returns us to a situation in which post–Part D induces rich distributional changes in real OOP but only a small additional mean shift conditional on time trend: with first-stage partial $F$-statistic is around 2. From the perspective of classical diagnostics, the instrument again looks fragile; from the perspective of distributional relevance, it remains informative.
Table 8 reports Q–LS and 2SLS estimates for this trend-adjusted real specification.
Table 8: Q–LS vs linear 2SLS: effect of real OOP risk on depression with time trend (2015\$)
(1) (2)
Q–LS 2SLS
Projected real OOP (Q–LS index, 2015 k\$) $0.07$
$(0.053)$
Real OOP spending (2015 k\$) $0.07$
$(0.079)$
Observations 16,295 16,295
Standard errors in parentheses.
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Notes: The dependent variable is an indicator for any depressive symptoms (CESD). Column (1) reports the Q–LS estimate obtained by regressing depression on the projected real OOP index $x_{\mathrm{QLS}}$ (in thousands of 2015 dollars), treating $x_{\mathrm{QLS}}$ as a generated regressor and using cluster-robust standard errors. Column (2) reports the linear 2SLS estimate where real OOP medical spending (in thousands of 2015 dollars) is instrumented with the post–Part D indicator. All specifications control for age, Medicaid coverage, Medicare coverage, sex, Hispanic origin, a linear time trend in years since 2000, and race (sracem), and include a constant. The sample consists of 16,295 person–year observations for HRS respondents aged 65 and older between 2000 and 2020, trimming the top 1% of the real OOP distribution. Standard errors are clustered at the individual level.
In this trend-adjusted real specification, Q–LS and 2SLS again deliver the same point estimate (0.07 per \$1,000 of real OOP). The main difference is precision: the Q–LS standard error is 0.05, whereas the 2SLS standard error is 0.07. Thus Q–LS reduces the confidence interval width by roughly one-third, but the estimate remains only marginally significant at best (p-value around 0.14). Economically, the estimates across the real-dollar specifications suggest a modest positive effect of OOP risk on depression, with the strongest evidence in the specification that treats Part D as a strong mean instrument and somewhat weaker evidence once secular trends are absorbed.
Taken together, the three designs – nominal weak-IV, real strong-IV, and real–with-trend -illustrate the two key empirical messages of the paper. First, the way in which a policy reform is summarized in the first stage (nominal vs real, with or without trend) can dramatically change whether the design looks weak or strong in mean terms, even when the underlying distributional shifts are similar. Second, when mean relevance is fragile but distributional relevance can be strong, Q–LS behaves like a natural complement to 2SLS: it recovers the same underlying signal where the design is well identified, and it delivers substantial precision gains relative to linear IV in settings where standard weak-IV diagnostics based on mean shifts understate the information content of the instrument.
8 Concluding remarks
Our results can be summarized along two dimensions. On the identification side, we formalize a notion of distributional relevance and show that even purely distributional instruments—for which $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ while $F_{X\mid Z}(\cdot\mid Z)$ varies nontrivially in ZZ—are sufficient to identify average structural functions in a nonseparable triangular model via the control function $V=F_{X\mid Z}(X\mid Z)$. This makes explicit how instruments that look weak in the mean can still be strong in the distribution, and connects the control–function identification argument to weak-IV concerns in linear models.
On the estimation side, we propose a simple, implementable estimator, Quantile Least Squares (Q–LS), which constructs an optimal distributional instrument from conditional quantiles and then uses it in a conventional linear IV step. In designs where policy reforms are primarily meant to reshape the distribution of risk rather than to move its mean, Q–LS complements the existing 2SLS and weak-IV toolkits by extracting a stronger effective first stage from the same policy variation, while leaving classical strong-IV applications essentially unchanged.
Q–LS exploits this structure by choosing, within a given dictionary of quantile-based functions, the instrument that best predicts $X$ in mean square error and then using it in a standard 2SLS procedure. We showed that Q–LS is consistent and asymptotically normal under distributional relevance, that it coincides with optimal 2SLS in Gaussian/location settings where mean relevance holds, and that it remains reliable when classical first-stage diagnostics indicate weak instruments from the perspective of the conditional mean. We also developed ridge and LASSO regularisation schemes that stabilise the first stage and mitigate many-instrument bias, while preserving the first-order asymptotics of Q–LS and providing practically useful variants (Q–LS–R, Q–LS–L1, Q–LS–L2) with distinct robustness properties.
Our Monte Carlo evidence highlights three main messages. First, when the instrument primarily shifts the tails of the $X$-distribution, Q–LS can recover the same limiting parameter as optimal 2SLS but with substantially smaller finite-sample variance, even in designs where the mean first stage is formally weak. Second, the regularised Q–LS variants are effective at controlling the dispersion of the estimator in many-instrument designs and under moderate misspecification of the quantile representation. Third, conventional IV estimators can behave erratically in these designs, with large RMSE and unstable confidence intervals, even when coverage remains nominal, whereas Q–LS delivers more concentrated sampling distributions.
The empirical application to Medicare Part D illustrates these features in a setting where policy reforms plausibly operate through the distribution of medical spending risk rather than its mean. In the nominal-dollar specifications, standard first-stage diagnostics classify the post–Part D indicator as a very weak instrument for mean OOP spending, and conventional 2SLS delivers large but imprecise estimates of the effect of OOP risk on depression. By contrast, Q–LS exploits the pronounced compression of the upper tail of the OOP distribution to construct a strong distributional first stage from the same policy variation, yielding substantially tighter confidence intervals while preserving the economic interpretation of the effect. When we deflate spending to 2015 dollars, the mean first stage becomes much stronger and conventional 2SLS performs well; in that case Q–LS and 2SLS again agree closely on the point estimates, but Q–LS offers at most modest efficiency gains. Taken together, the simulations and the Medicare Part D evidence suggest viewing Q–LS as a complement to standard IV: it nests the optimal mean-based instrument in classical designs, but remains informative in applications where instruments are weak in the mean and strong in the distribution.
The distributional relevance framework extends beyond linear IV. In an appendix, we show that purely distributional instruments can identify average structural functions in nonseparable triangular models via a control-function representation based on the conditional distribution $F_{X}(X\mid Z)$. This connects our approach to the literature on nonparametric IV and completeness, and suggests several directions for future work: deriving more primitive conditions that guarantee nondegenerate optimal instruments; developing formal diagnostics for the linear quantile representation and for instrument-instability; designing data-driven procedures for selecting quantile grids and penalties; and applying Q–LS in empirical settings where policy or institutional variation plausibly operates through changes in risk exposure rather than simple mean shifts.
Appendix A Nonseparable triangular models and distributional relevance
This appendix collects the extension of distributional relevance and the control-function approach to nonseparable triangular models, and outlines a semiparametric estimator for the average structural function. The material follows the draft text in Sections AA.5 of the earlier version of the paper, with notational adjustments for consistency.
We build on the nonseparable triangular framework of Chernozhukov et al. (2020) to formalize a notion of distributional relevance of instruments. Our goal is to show that identification of the average structural function can be achieved even when instruments are “weak” in the classical mean sense, provided they are sufficiently strong in a distributional sense.
A.1 Triangular model and object of interest
Let $(Y,X,Z)$ be random variables defined on a common probability space. We consider the following nonseparable triangular system:
$Y$ $=f(X,Z_{1},\varepsilon),$ (26)
$X$ $=g(Z,\eta),$ (27)
where $f$ and $g$ are unknown measurable functions, and $\varepsilon,\eta$ are unobserved disturbances. The variable $Z=(Z_{1},Z_{2})^{\prime}$ is a vector of instruments.
Assumption A.1 (Exogeneity and normalization).
The pair $(\varepsilon,\eta)$ is independent of $Z$, and $\eta\sim U(0,1)$.
We are interested in the average structural function (ASF)
$\mu(x,z_{1}):=\mathbb{E}\big[f(x,z_{1},\varepsilon)\big],$ (28)
which captures the mean potential outcome if the endogenous and exogenous regressors were set to the value $x$ and $z_{1}$.
A.2 Control-function representation via the conditional distribution of $X$
Let $F_{X}(x\mid z)$ denote the conditional distribution function of $X$ given $Z=z$:
$F_{X}(x\mid z):=\mathbb{P}(X\leq x\mid Z=z).$
Define the control variable
$V:=F_{X}(X\mid Z).$ (29)
Following Chernozhukov et al. (2020), we impose a monotonicity condition on the first stage.
Assumption A.2 (Monotone first stage).
For each $z$ in the support of $Z$, the map $\eta\mapsto g(z,\eta)$ in (27) is strictly increasing and continuous.
Under Assumptions A.1A.2, the first stage admits a conditional quantile representation; see Chernozhukov et al. (2020):
$X=g(Z,\eta)\quad\Longrightarrow\quad F_{X}(X\mid Z)=\eta\quad\text{a.s.,}$
so that
$V=F_{X}(X\mid Z)=\eta\quad\text{a.s.}$ (30)
In particular, $V$ is independent of $Z$ and can be used as a control function for the endogeneity of $X$.
A key implication of (30) and Assumption A.1 is that $\varepsilon$ is independent of $(X,Z)$ conditional on $V$:
$\varepsilon\perp(X,Z)\mid V.$ (31)
It follows from (26) and (31) that
$\mathbb{E}[Y\mid X=x,Z_{1}=z_{1},V=v]=\mathbb{E}[f(x,z_{1},\varepsilon)\mid V=v].$ (32)
To recover the ASF $\mu(x,z_{1})$ in (28), we rely on a support condition that is standard in control-function identification results.
Assumption A.3 (Support / overlap).
For each $x$ in the interior of the support of $X$, the conditional support of $V$ given $X=x$ coincides with the marginal support of $V$:
$\operatorname{supp}(V\mid X=x)=\operatorname{supp}(V).$
Under Assumptions A.1A.3, one obtains the usual control-function representation for the ASF:
$\mu(x,z_{1})=\mathbb{E}[f(x,z_{1},\varepsilon)]=\int\mathbb{E}[Y\mid X=x,Z_{1}=z_{1},V=v]\,dF_{V}(v).$ (33)
A.3 Mean relevance and distributional relevance in an $L^{2}$ space
The notions of mean relevance and distributional relevance in Definitions 1 and 2 extend verbatim to this nonparametric setting: we view $m(z)=\mathbb{E}[X\mid Z=z]$ as an element of $L^{2}(P_{Z})$, and $F_{X}(\cdot\mid z)$ as an element of $L^{2}(\lambda)$. We then define purely distributional instruments as in Definition 3.
A.4 Identification of the ASF under purely distributional relevance
We now show that the ASF in (28) remains nonparametrically identified under purely distributional relevance, even though the instruments are completely irrelevant for the conditional mean of $X$.
Theorem 1 (ASF identification with purely distributional instruments).
Suppose Assumptions A.1, A.2, and A.3 hold. Assume further that $Z$ includes only purely distributional instruments for $X$ in the sense of Definition 3. Then, for every $x$ in the interior of the support of $X$, the ASF $\mu(x,z_{1})=\mathbb{E}[f(x,z_{1},\varepsilon)]$ is nonparametrically identified and admits the representation
$\mu(x,z_{1})=\int\mathbb{E}[Y\mid X=x,Z_{1}=z_{1},V=v]\,dF_{V}(v).$ (34)
In particular, identification of $\mu(x,z_{1})$ does not require any mean relevance of the instruments; it is enough that $Z$ is distributionally relevant for $X$.
Proof.
By Assumptions A.1 and A.2, we have $V=F_{X}(X\mid Z)=\eta$ almost surely, and $V$ is independent of $Z$. Moreover, the joint system (26)–(27) with $(\varepsilon,\eta)\perp Z$ implies that $\varepsilon$ is independent of $(X,Z)$ conditional on $V$, as in (31). Therefore,
$\mathbb{E}[Y\mid X=x,Z_{1}=z_{1},V=v]=\mathbb{E}[f(x,z_{1},\varepsilon)\mid V=v],$
as in (32). By Assumption A.3, the conditional support of $V$ given $X=x$ coincides with its marginal support. Using the law of total expectation, we can write
$\mu(x,z_{1})=\mathbb{E}[f(x,z_{1},\varepsilon)]=\int\mathbb{E}[f(x,z_{1},\varepsilon)\mid V=v]\,dF_{V}(v)=\int\mathbb{E}[Y\mid X=x,Z_{1}=z_{1},V=v]\,dF_{V}(v),$
which yields (34). The purely distributional instrument condition ensures that $F_{X}(\cdot\mid Z)$ varies with $Z$ in an $L^{2}(\lambda)$ sense, so that the control variable $V=F_{X}(X\mid Z)$ is non-degenerate. The lack of mean relevance, $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$, does not affect this argument, since the control-function approach does not require $Z$ to move the mean of $X$. ∎
A.5 Semiparametric estimation with distributionally relevant instruments
We briefly outline a semiparametric estimator of the ASF based on a two-step plug-in approach: (i) estimate the control function $V=F_{X}(X\mid Z)$, and (ii) estimate the conditional mean function $m(x,z_{1},v):=\mathbb{E}[Y\mid X=x,Z_{1}=z_{1},V=v]$ using flexible basis expansions. The estimator of the ASF is then obtained by plugging in the estimated control function and conditional mean.
Let $\{(Y_{i},X_{i},Z_{i})\}_{i=1}^{n}$ be an i.i.d. sample. The estimator proceeds as follows.
  1. Step 1.
    Estimate the conditional distribution (or quantile process) of $X$ given $Z$ and construct an estimate $\hat{V}_{i}$ of the control variable $V_{i}=F_{X}(X_{i}\mid Z_{i})$.
  2. Step 2.
[…]
    $\mu(x,z_{1})=\int m(x,z_{1},v)\,dF_{V}(v)$
    to construct a plug-in estimator $\hat{\mu}(x,z_{1})$.
The first stage can be implemented using series quantile regression as in Chernozhukov et al. (2015) and Chernozhukov et al. (2020). A finite grid of quantile indices $\mathcal{T}_{n}=\{\tau_{1},\dots,\tau_{K_{n}}\}$ is used to estimate conditional quantiles $Q_{X}(\tau_{k}\mid Z)$ by series quantile regression in $Z$, and the conditional CDF at $(X_{i},Z_{i})$ is approximated by
$\hat{F}_{X}(X_{i}\mid Z_{i}):=\mathbbm{1}\big\{\hat{Q}_{X}(\tau_{k}\mid Z_{i})\leq X_{i}\big\}.$ (35)
The resulting estimated control variable is $\hat{V}_{i}=\hat{F}_{X}(X_{i}\mid Z_{i})$.
In the second stage, one specifies a basis $w(X,Z_{1},V)$ in $(X,Z_{1},V)$ and estimates
$m(x,z_{1},v)\approx w(x,z_{1},v)^{\prime}\beta$
by least squares, replacing $V_{i}$ by $\hat{V}_{i}$. The ASF is then estimated by
$\hat{\mu}(x,z_{1}):=\frac{1}{n}\sum_{i=1}^{n}\hat{m}(x,z_{1},\hat{V}_{i})=\frac{1}{n}\sum_{i=1}^{n}w(x,z_{1},\hat{V}_{i})^{\prime}\hat{\beta}$. (36)
Under standard smoothness and sieve-approximation assumptions on $F_{X}(x\mid z)$ and $m(x,z_{1},v)$, and suitable growth rates for the sieve dimensions and the number of quantiles $K_{n}$, one can show that $\hat{\mu}(x,z_{1})$ is consistent and asymptotically normal, with an influence function that is orthogonal to the first-stage estimation error in $\hat{V}$; see Chernozhukov et al. (2020) for related results.
Appendix B Detailed Proofs of Lemma and Propositions
This appendix collects the main proofs and clarifies the role of the Q–LS instrument as an $L^{2}$ projection. Throughout we use the same notation and assumptions as in the main text and treat the number of quantile basis functions $J$ as fixed (or growing slowly enough with $n$ that standard sieve arguments apply).
B.1 Setup and assumptions
Let $\{(Y_{i},X_{i},Z_{i})\}_{i=1}^{n}$ be i.i.d. observations, where
$Z_{i}=\begin{pmatrix}Z_{1i}\\ Z_{2i}\end{pmatrix}$,
$Z_{1i}\in\mathbb{R}^{d_{1}}$ denotes the included exogenous controls, and $Z_{2i}$ is the policy component (e.g. the post–Part D indicator). The structural equation is
$Y_{i}=\beta_{0}X_{i}+Z_{1i}^{\prime}\gamma_{0}+\varepsilon_{i},\qquad\mathbb{E}[\varepsilon_{i}\mid Z_{i}]=0.$ (37)
We stack the structural regressors as
$S_{0}(Z_{i}):=\begin{pmatrix}X_{i}\\ Z_{1i}\\ 1\end{pmatrix}\in\mathbb{R}^{d_{s}},$
and write $S_{i}:=S_{0}(Z_{i})$ for short.
The Q–LS instrument is built from a finite-dimensional dictionary of transformations of $Z_{i}$. Let $G_{n}(Z_{i})\in\mathbb{R}^{J_{n}}$ denote the $J_{n}$-dimensional dictionary at sample size $n$. In applications, $G_{n}(Z_{i})$ collects (functions of) conditional quantiles of $X_{i}$ given $Z_{i}$ evaluated on a finite grid $\mathcal{T}_{n}$ of quantile indices; the dimension $J_{n}$ is allowed to depend on $|\mathcal{T}_{n}|$.
We recall and slightly specialize the assumptions from the main text.
Assumption B.1 (Sampling and moments).
(i) $\{(Y_{i},X_{i},Z_{i})\}_{i=1}^{n}$ are i.i.d. and $\mathbb{E}\|S_{0}(Z_{i})\|^{4}<\infty$, $\mathbb{E}\|G_{n}(Z_{i})\|^{4}<\infty$ for all $n$.
(ii) The matrices
$A_{n}:=\mathbb{E}\big[S_{0}(Z_{i})X_{i}\big],\qquad B_{n}:=\mathbb{E}\big[S_{0}(Z_{i})G_{n}(Z_{i})^{\prime}\big]$
are finite and have full column rank uniformly in $n$.
(iii) The number of dictionary terms $J_{n}$ (and, equivalently, the quantile grid size $|\mathcal{T}_{n}|$) satisfies
$J_{n}\to\infty\quad\text{and}\quad\frac{J_{n}^{2}}{n}\to 0\quad\text{as }n\to\infty.$
Assumption B.2 (Distributional relevance).
Let $\mathcal{H}_{n}$ denote the closed linear span in $L^{2}$ of the dictionary $\{G_{n}(Z_{i})\}$ and the included controls $Z_{1i}$ and a constant:
$\mathcal{H}_{n}:=\overline{\text{span}\big\{G_{n}(Z_{i}),\,Z_{1i},\,1\big\}}\subset L^{2}.$
There exists a sequence $h_{n}\in\mathcal{H}_{n}$ such that
$Cov(X_{i},h_{n}(Z_{i}))\neq 0\quad\text{for all large }n.$
Assumption B.2 is the sieve version of the distributional relevance Assumption 2 in the main text: at least one element of the instrument space is correlated with $X_{i}$, even if the conditional mean $\mathbb{E}[X_{i}\mid Z_{i}]$ is flat in $Z_{2i}$.
We define the population $L^{2}$-projection of $X_{i}$ onto the infinite-dimensional limit space
$\mathcal{H}_{\infty}:=\overline{\bigcup_{n\geq 1}\mathcal{H}_{n}}$
by
$h^{\star}\in\arg\min_{h\in\mathcal{H}_{\infty}}\mathbb{E}\big[(X_{i}-h(Z_{i}))^{2}\big],$ (38)
and denote the corresponding projection error by
$r^{\star}_{i}:=X_{i}-h^{\star}(Z_{i}).$
Assumption B.3 (Approximation of the optimal instrument).
For each $n$, define
$h_{n}(Z_{i}):=G_{n}(Z_{i})^{\prime}w_{0n},\qquad w_{0n}\in\arg\min_{w\in\mathbb{R}^{J_{n}}}\mathbb{E}\big[(h^{\star}(Z_{i})-G_{n}(Z_{i})^{\prime}w)^{2}\big].$
Then the sieve approximates $h^{\star}$ in $L^{2}$:
$\mathbb{E}\big[(h^{\star}(Z_{i})-h_{n}(Z_{i}))^{2}\big]\to 0\quad\text{as }n\to\infty.$
Assumption B.3 is the usual sieve approximation condition: as the dictionary grows (via a finer quantile grid and more basis terms), linear combinations of $G_{n}(Z_{i})$ approximate the optimal projection $h^{\star}$ in mean square.
B.2 Instrument relevance of the optimal projection
We first show that the $L^{2}$-projection $h^{\star}$ is a relevant instrument whenever the dictionary is distributionally relevant.
Lemma B.1.
Let $h^{\star}$ be defined as in (38). Then either
  1. (a)
    $h^{\star}(Z_{i})=0$ almost surely, or
  2. (b)
    $\mathbb{E}[X_{i}h^{\star}(Z_{i})]=\mathbb{E}[(h^{\star}(Z_{i}))^{2}]>0$.
Under Assumption B.2, case (b) must hold, so $Cov(X_{i},h^{\star}(Z_{i}))\neq 0$.
Proof.
By the $L^{2}$-projection property, the error $r^{\star}_{i}=X_{i}-h^{\star}(Z_{i})$ is orthogonal to the space $\mathcal{H}_{\infty}$:
$\mathbb{E}[r^{\star}_{i}h(Z_{i})]=0\quad\text{for all }h\in\mathcal{H}_{\infty}.$
In particular, taking $h=h^{\star}$,
$\mathbb{E}[r^{\star}_{i}h^{\star}(Z_{i})]=0.$
Thus
$\mathbb{E}[X_{i}h^{\star}(Z_{i})]$ $=\mathbb{E}[(h^{\star}(Z_{i})+r^{\star}_{i})h^{\star}(Z_{i})]$
$=\mathbb{E}[(h^{\star}(Z_{i}))^{2}]+\mathbb{E}[r^{\star}_{i}h^{\star}(Z_{i})]$
$=\mathbb{E}[(h^{\star}(Z_{i}))^{2}]\geq 0.$
If $h^{\star}(Z_{i})=0$ almost surely, then $\mathbb{E}[X_{i}h^{\star}(Z_{i})]=0$. If $\P(|h^{\star}(Z_{i})|>0)>0$, then $\mathbb{E}[(h^{\star}(Z_{i}))^{2}]>0$, hence $\mathbb{E}[X_{i}h^{\star}(Z_{i})]>0$.
Now suppose Assumption B.2 holds. For each $n$, write
$X_{i}=h^{\star}(Z_{i})+r^{\star}_{i},\qquad r^{\star}_{i}\perp\mathcal{H}_{\infty}.$
For any $h_{n}\in\mathcal{H}_{n}\subset\mathcal{H}_{\infty}$,
$Cov(X_{i},h_{n}(Z_{i}))=Cov(h^{\star}(Z_{i}),h_{n}(Z_{i}))+Cov(r^{\star}_{i},h_{n}(Z_{i}))=Cov(h^{\star}(Z_{i}),h_{n}(Z_{i})),$
because $r^{\star}_{i}$ is orthogonal to $\mathcal{H}_{\infty}$. If $h^{\star}(Z_{i})=0$ a.s., then $Cov(X_{i},h_{n}(Z_{i}))=0$ for all $h_{n}$, which contradicts Assumption B.2. Hence $h^{\star}$ cannot be identically zero, so we are in case (b) and $Cov(X_{i},h^{\star}(Z_{i}))=\mathbb{E}[(h^{\star}(Z_{i}))^{2}]>0$. ∎
Remark 1.
This is a direct application of the orthogonal projection theorem in Hilbert spaces: for any closed subspace $\mathcal{M}$ of $L^{2}$, the projection $P_{\mathcal{M}}X$ satisfies $\mathbb{E}[(X-P_{\mathcal{M}}X)Z]=0$ for all $Z\in\mathcal{M}$, so $\mathbb{E}[XZ]=\mathbb{E}[(P_{\mathcal{M}}X)Z]$ (e.g. Kreyszig 1978, Ch. 3).
B.3 Equivalence of 2SLS and OLS with projected regressors
We next recall the standard algebra showing that 2SLS with instruments $M_{i}$ is equivalent to OLS on the regressors projected onto the instrument space.
Let
$S_{i}=\begin{pmatrix}X_{i}\\ Z_{1i}\\ 1\end{pmatrix},\qquad M_{i}=M(Z_{i})\in\mathbb{R}^{L}$
denote the stacked structural regressors and instruments, and define the $n\times d_{s}$ and $n\times L$ matrices
$\mathbf{S}=(S_{1}^{\prime},\dots,S_{n}^{\prime})^{\prime},\qquad\mathbf{M}=(M_{1}^{\prime},\dots,M_{n}^{\prime})^{\prime}.$
Also set $\mathbf{Y}=(Y_{1},\dots,Y_{n})^{\prime}$.
The 2SLS estimator using instruments $\mathbf{M}$ is
$\hat{\theta}^{\mathrm{2SLS}}=(\mathbf{S}^{\prime}\mathbf{P}_{M}\mathbf{S})^{-1}\mathbf{S}^{\prime}\mathbf{P}_{M}\mathbf{Y},\qquad\mathbf{P}_{M}:=\mathbf{M}(\mathbf{M}^{\prime}\mathbf{M})^{-1}\mathbf{M}^{\prime}.$ (39)
Let $\tilde{\mathbf{S}}:=\mathbf{P}_{M}\mathbf{S}$ denote the sample projection of the regressors onto the instrument space; its $i$th row is $\tilde{S}_{i}:=\mathbb{E}_{n}[S_{i}\mid M_{i}]$. Consider the OLS regression
$\mathbf{Y}=\tilde{\mathbf{S}}\,\tilde{\theta}+u.$
Lemma B.2.
Suppose $\mathbf{M}$ has full column rank and $\mathbf{S}^{\prime}\mathbf{P}_{M}\mathbf{S}$ is nonsingular. Then the 2SLS estimator $\hat{\theta}^{\mathrm{2SLS}}$ in (39) coincides with the OLS estimator $\tilde{\theta}$ from the regression of $\mathbf{Y}$ on $\tilde{\mathbf{S}}$. In particular, the 2SLS coefficient on $X_{i}$ equals the OLS coefficient on its instrumented version $\tilde{X}_{i}$.
Proof.
By definition,
$\tilde{\mathbf{S}}=\mathbf{P}_{M}\mathbf{S}=\mathbf{M}(\mathbf{M}^{\prime}\mathbf{M})^{-1}\mathbf{M}^{\prime}\mathbf{S}.$
The OLS estimator from $\mathbf{Y}=\tilde{\mathbf{S}}\,\tilde{\theta}+u$ is
$\tilde{\theta}$ $=(\tilde{\mathbf{S}}^{\prime}\tilde{\mathbf{S}})^{-1}\tilde{\mathbf{S}}^{\prime}\mathbf{Y}$
$=(\mathbf{S}^{\prime}\mathbf{P}_{M}^{\prime}\mathbf{P}_{M}\mathbf{S})^{-1}\mathbf{S}^{\prime}\mathbf{P}_{M}^{\prime}\mathbf{Y}.$
Since $\mathbf{P}_{M}$ is a symmetric idempotent projection, $\mathbf{P}_{M}^{\prime}=\mathbf{P}_{M}$ and $\mathbf{P}_{M}^{2}=\mathbf{P}_{M}$, so
$\tilde{\theta}=(\mathbf{S}^{\prime}\mathbf{P}_{M}\mathbf{S})^{-1}\mathbf{S}^{\prime}\mathbf{P}_{M}\mathbf{Y}=\hat{\theta}^{\mathrm{2SLS}}.$
The statement about the coefficient on $X_{i}$ follows by partitioning $\theta=(\beta,\gamma^{\prime},\alpha)^{\prime}$ corresponding to $(X_{i},Z_{1i},1)$. ∎
Remark 2.
This is the usual equivalence between 2SLS and OLS on the fitted values of the endogenous regressors; see, e.g., Wooldridge (2010) and Angrist and Imbens (1995).
B.4 Definition and consistency of the Q–LS estimator
In the Q–LS procedure we choose a weight vector $w_{n}\in\mathbb{R}^{J_{n}}$ that makes the projected instrument $G_{n}(Z_{i})^{\prime}w_{n}$ as close as possible, in moment terms, to the optimal projection $h^{\star}(Z_{i})$.
Define the sample criterion
$\hat{Q}_{n}(w):=\Big\|\frac{1}{n}\sum_{i=1}^{n}S_{i}\big(X_{i}-G_{n}(Z_{i})^{\prime}w\big)\Big\|^{2},\qquad w\in\mathbb{R}^{J_{n}},$ (40)
and let
$\hat{w}_{n}\in\arg\min_{w\in\mathbb{R}^{J_{n}}}\hat{Q}_{n}(w).$
The Q–LS instrument is then
$\hat{h}_{n}(Z_{i}):=G_{n}(Z_{i})^{\prime}\hat{w}_{n},$
and the Q–LS IV estimator $\hat{\beta}_{n}$ is defined as the 2SLS coefficient on $X_{i}$ in (37) using $\hat{h}_{n}(Z_{i})$ as the excluded instrument and $Z_{1i}$ as included controls (equivalently, OLS on the projected regressor by Lemma B.2).
The corresponding population criterion is
$Q_{n}(w):=\Big\|\mathbb{E}\big[S_{0}(Z_{i})\big(X_{i}-G_{n}(Z_{i})^{\prime}w\big)\big]\Big\|^{2},$ (41)
and we denote by $w_{0n}$ a minimizer of $Q_{n}(w)$ (which coincides with the $L^{2}$ projection of $X_{i}$ onto the span of $G_{n}(Z_{i})$, as “seen through” the moments with $S_{0}(Z_{i})$).
Proposition B.1 (Consistency of Q–LS).
Under Assumptions B.1B.3 and distributional relevance (Assumption B.2), if $Q_{n}(w)$ has a unique minimizer $w_{0n}$, then
$\hat{w}_{n}\overset{p}{\longrightarrow}w_{0n},\qquad\hat{h}_{n}(Z_{i})=G_{n}(Z_{i})^{\prime}\hat{w}_{n}\overset{L^{2}}{\longrightarrow}h^{\star}(Z_{i}),$
and the Q–LS estimator $\hat{\beta}_{n}$ is consistent:
$\hat{\beta}_{n}\overset{p}{\longrightarrow}\beta_{0}.$
Proof.
Step 1: Uniform convergence of the criterion. For fixed $n$, define
$m_{n}(w):=\mathbb{E}\big[S_{0}(Z_{i})\big(X_{i}-G_{n}(Z_{i})^{\prime}w\big)\big],$
so that $Q_{n}(w)=\|m_{n}(w)\|^{2}$. The sample analogue is
$\hat{m}_{n}(w):=\frac{1}{n}\sum_{i=1}^{n}S_{i}\big(X_{i}-G_{n}(Z_{i})^{\prime}w\big),\qquad\hat{Q}_{n}(w)=\|\hat{m}_{n}(w)\|^{2}.$
Because $G_{n}(Z_{i})$ is finite-dimensional and has bounded fourth moments (Assumption B.1), $\hat{m}_{n}(w)$ is a finite-dimensional empirical average with linear dependence on $w$. Standard multivariate LLN and a uniform LLN for parametric families (e.g. Newey and McFadden 1994, Sec. 2.4) imply that on any compact set $\mathcal{W}_{n}\subset\mathbb{R}^{J_{n}}$,
$\sup_{w\in\mathcal{W}_{n}}\big\|\hat{m}_{n}(w)-m_{n}(w)\big\|\overset{p}{\longrightarrow}0,$
hence also
$\sup_{w\in\mathcal{W}_{n}}\big|\hat{Q}_{n}(w)-Q_{n}(w)\big|\overset{p}{\longrightarrow}0,$
using the Lipschitz continuity of $a\mapsto\|a\|^{2}$.
Step 2: Consistency of $\hat{w}_{n}$. By Assumption B.1(ii) and the quadratic form of $Q_{n}(w)$, the minimizer $w_{0n}$ is unique and lies in a bounded region of $\mathbb{R}^{J_{n}}$. Choosing $\mathcal{W}_{n}$ large enough to contain $w_{0n}$ with probability approaching one, the usual M–estimation argument (Newey and McFadden 1994, Thm. 2.1) yields
$\hat{w}_{n}\overset{p}{\longrightarrow}w_{0n}.$
Step 3: Approximation of $h^{\star}$. By Assumption B.3,
$\mathbb{E}\big[(h^{\star}(Z_{i})-h_{n}(Z_{i}))^{2}\big]=\mathbb{E}\big[(h^{\star}(Z_{i})-G_{n}(Z_{i})^{\prime}w_{0n})^{2}\big]\to 0.$
Moreover, the convergence $\hat{w}_{n}\to_{p}w_{0n}$ and bounded second moments of $G_{n}(Z_{i})$ imply
$\mathbb{E}\big[(\hat{h}_{n}(Z_{i})-h_{n}(Z_{i}))^{2}\big]=\mathbb{E}\big[(G_{n}(Z_{i})^{\prime}(\hat{w}_{n}-w_{0n}))^{2}\big]=o_{p}(1).$
Combining the two displays gives
$\mathbb{E}\big[(\hat{h}_{n}(Z_{i})-h^{\star}(Z_{i}))^{2}\big]\to_{p}0,$
that is, $\hat{h}_{n}(Z_{i})\overset{L^{2}}{\to}h^{\star}(Z_{i})$.
Step 4: Consistency of $\hat{\beta}_{n}$. Consider the “oracle” IV estimator that uses the optimal instrument $h^{\star}(Z_{i})$:
$\beta_{n}^{\star}:=\frac{\frac{1}{n}\sum_{i=1}^{n}h^{\star}(Z_{i})\big(Y_{i}-Z_{1i}^{\prime}\gamma_{0}\big)}{\frac{1}{n}\sum_{i=1}^{n}h^{\star}(Z_{i})X_{i}}.$
By the structural equation (37) and $\mathbb{E}[\varepsilon_{i}\mid Z_{i}]=0$,
$\mathbb{E}\big[h^{\star}(Z_{i})(Y_{i}-Z_{1i}^{\prime}\gamma_{0})\big]=\beta_{0}\,\mathbb{E}\big[h^{\star}(Z_{i})X_{i}\big].$
Lemma B.1 and distributional relevance ensure $\mathbb{E}[h^{\star}(Z_{i})X_{i}]\neq 0$, so by the LLN, $\beta_{n}^{\star}\to_{p}\beta_{0}$.
The actual Q–LS estimator replaces $h^{\star}(Z_{i})$ by $\hat{h}_{n}(Z_{i})$ in the same IV formula. Because $\hat{h}_{n}(Z_{i})\to h^{\star}(Z_{i})$ in $L^{2}$ and is uniformly square-integrable, the difference between the sample moments with $\hat{h}_{n}$ and with $h^{\star}$ is $o_{p}(1)$ (see, e.g., Newey and McFadden 1994, Sec. 6.2, on IV with generated regressors). Therefore, the Q–LS IV estimator $\hat{\beta}_{n}$ has the same probability limit as $\beta_{n}^{\star}$, namely $\beta_{0}$. ∎
B.5 Asymptotic normality of Q–LS
We now derive the asymptotic distribution of $\hat{\beta}_{n}$ under a fixed dictionary size $J_{n}=J$; extending to slowly growing $J_{n}$ requires additional notation but follows the same steps.
Define
$\phi_{i}(\beta,h):=h(Z_{i})\big(Y_{i}-\beta X_{i}-Z_{1i}^{\prime}\gamma_{0}\big),$
and consider the IV moment condition
$\mathbb{E}[\phi_{i}(\beta_{0},h^{\star})]=0.$
The Q–LS estimator $\hat{\beta}_{n}$ solves the sample moment equation
$\hat{m}_{\beta}(\beta,\hat{h}_{n}):=\frac{1}{n}\sum_{i=1}^{n}\hat{h}_{n}(Z_{i})\big(Y_{i}-\beta X_{i}-Z_{1i}^{\prime}\hat{\gamma}_{n}\big)=0,$
where $\hat{\gamma}_{n}$ denotes the companion coefficient on $Z_{1i}$ in the same IV regression. For clarity we suppress $\gamma$ in the notation below and focus on the scalar coefficient $\beta$.
Let
$A_{0}:=\mathbb{E}[h^{\star}(Z_{i})X_{i}],\qquad\Omega_{0}:=\mathbb{E}[(h^{\star}(Z_{i})\varepsilon_{i})^{2}].$
Proposition B.2 (Asymptotic normality of Q–LS).
Under Assumptions B.1B.3, distributional relevance, and fixed $J_{n}=J$, the Q–LS estimator satisfies
$\sqrt{n}(\hat{\beta}_{n}-\beta_{0})\;\overset{d}{\longrightarrow}\;N\big(0,V_{\mathrm{QLS}}\big),\qquad V_{\mathrm{QLS}}:=A_{0}^{-2}\Omega_{0}.$
In the stacked system including $\gamma_{0}$, this corresponds to the first diagonal element of the usual IV sandwich covariance matrix.
Proof.
Step 1: Linearization of the moment condition. By definition,
$0=\hat{m}_{\beta}(\hat{\beta}_{n},\hat{h}_{n})=\hat{m}_{\beta}(\beta_{0},\hat{h}_{n})+\left.\frac{\partial\hat{m}_{\beta}(\beta,\hat{h}_{n})}{\partial\beta}\right|_{\beta=\beta_{0}}(\hat{\beta}_{n}-\beta_{0})+o_{p}(n^{-1/2}).$
We can write
$\hat{m}_{\beta}(\beta_{0},\hat{h}_{n})=\frac{1}{n}\sum_{i=1}^{n}\hat{h}_{n}(Z_{i})\varepsilon_{i}+o_{p}(n^{-1/2}),$
where the $o_{p}(n^{-1/2})$ term collects the contribution of estimating $\gamma_{0}$; standard IV theory implies that this contribution does not affect the first-order distribution of $\hat{\beta}_{n}$ (see, e.g., Wooldridge 2010, ch. 5).
Under $\hat{h}_{n}(Z_{i})\to h^{\star}(Z_{i})$ in $L^{2}$ and a Lindeberg condition on $h^{\star}(Z_{i})\varepsilon_{i}$, the CLT gives
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{h}_{n}(Z_{i})\varepsilon_{i}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h^{\star}(Z_{i})\varepsilon_{i}+o_{p}(1)\;\overset{d}{\longrightarrow}\;N(0,\Omega_{0}).$
Step 2: Limit of the derivative. The derivative of $\hat{m}_{\beta}(\beta,\hat{h}_{n})$ with respect to $\beta$ is
$\frac{\partial\hat{m}_{\beta}(\beta,\hat{h}_{n})}{\partial\beta}=-\frac{1}{n}\sum_{i=1}^{n}\hat{h}_{n}(Z_{i})X_{i}.$
By the LLN and $\hat{h}_{n}(Z_{i})\to h^{\star}(Z_{i})$ in $L^{2}$,
$\frac{\partial\hat{m}_{\beta}(\beta_{0},\hat{h}_{n})}{\partial\beta}\;\overset{p}{\longrightarrow}\;A_{0}=\mathbb{E}[h^{\star}(Z_{i})X_{i}],$
which is nonzero by Lemma B.1.
Step 3: Putting pieces together. Combining the linearization and the previous limits,
$\sqrt{n}(\hat{\beta}_{n}-\beta_{0})=A_{0}^{-1}\,\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h^{\star}(Z_{i})\varepsilon_{i}+o_{p}(1)\;\overset{d}{\longrightarrow}\;N(0,A_{0}^{-2}\Omega_{0}).$
In the finite-dimensional case considered here, this variance coincides with the standard 2SLS variance formula evaluated at the optimal instrument $h^{\star}(Z_{i})$. In practice, we estimate the asymptotic variance with the usual cluster-robust IV sandwich estimator, replacing $h^{\star}$ by $\hat{h}_{n}$ and sample expectations by empirical averages.
B.6 Equivalence with linear 2SLS under mean-shift designs
Finally, we formalize the statement that when the first stage is exactly linear in the original policy instrument and controls, and the Q–LS dictionary contains this linear space, Q–LS coincides with ordinary 2SLS.
Suppose:
  1. (i)
    The dictionary includes the original scalar instrument $Z_{2i}$, the controls $Z_{1i}$, and a constant, so that
    $\mathcal{H}_{n}\supset\text{span}\{Z_{2i},Z_{1i},1\}.$
  2. (ii)
    The true first stage is linear:
    $X_{i}=\pi_{0}Z_{2i}+Z_{1i}^{\prime}\pi_{1}+u_{i},\qquad\mathbb{E}[u_{i}\mid Z_{i}]=0.$ (42)
Proposition B.3 (Equivalence with linear 2SLS).
Under (i)–(ii),
  1. (a)
    The $L^{2}$-projection $h^{\star}(Z_{i})$ of $X_{i}$ onto $\mathcal{H}_{\infty}$ is exactly the linear conditional mean
    $h^{\star}(Z_{i})=\mathbb{E}[X_{i}\mid Z_{1i},Z_{2i}]=\pi_{0}Z_{2i}+Z_{1i}^{\prime}\pi_{1}+\alpha_{0}$
    for some constant $\alpha_{0}$.
  2. (b)
    The Q–LS estimator $\hat{\beta}_{n}$ is numerically identical to the standard 2SLS estimator that uses $Z_{2i}$ as the excluded instrument and $Z_{1i}$ as controls.
Proof.
(a) Because $\mathcal{H}_{\infty}$ contains all linear combinations of $(Z_{2i},Z_{1i},1)$, the best $L^{2}$ approximation to $X_{i}$ in $\mathcal{H}_{\infty}$ under (42) is its linear conditional mean
$\mathbb{E}[X_{i}\mid Z_{1i},Z_{2i}]=\pi_{0}Z_{2i}+Z_{1i}^{\prime}\pi_{1}+\alpha_{0}.$
(b) In this case, the population Q–LS weights recover exactly the linear first-stage coefficients, and the fitted values $\hat{h}_{n}(Z_{i})$ coincide (up to sampling noise) with the usual linear first-stage fitted values from regressing $X_{i}$ on $(Z_{2i},Z_{1i})$. By Lemma B.2, 2SLS with instrument vector $(Z_{2i},Z_{1i},1)$ and regressor vector $(X_{i},Z_{1i},1)$ is equivalent to OLS of $Y_{i}$ on the projected regressors $(\tilde{X}_{i},\tilde{Z}_{1i},1)$, where $\tilde{X}_{i}=\mathbb{E}[X_{i}\mid Z_{1i},Z_{2i}]$ and $\tilde{Z}_{1i}=Z_{1i}$.
The Q–LS estimator uses $\hat{h}_{n}(Z_{i})$ in place of $\tilde{X}_{i}$ but, under an exactly linear first stage and a dictionary that contains the linear space, $\hat{h}_{n}(Z_{i})$ equals $\tilde{X}_{i}$ in finite samples up to numerical precision. Therefore, Q–LS reproduces the usual 2SLS coefficient on $X_{i}$. ∎
Remark 3.
Proposition B.3 formalizes the idea that when the instrument only induces a mean shift in $X_{i}$ and the first stage is linear, Q–LS does not change the linear IV estimate. The main gains of Q–LS arise precisely when instruments are distributionally relevant but mean–weak in the sense that $\operatorname{Var}(\mathbb{E}[X_{i}\mid Z_{i}])$ is small even though $F_{X\mid Z}(\cdot\mid Z_{i})$ varies substantially in $Z_{i}$.
Table 9: Selected evidence on mean spending vs. financial-risk protection
Study Setting / population Mean-level outcome Risk / tail outcome
Finkelstein and McKnight (2008) Introduction of Medicare (1965), U.S. adults 65+ No statistically significant effect on out-of-pocket (OOP) spending for the bottom 75% of the pre-Medicare distribution. Top quartile of OOP spending for the elderly falls by about 40%; OOP spending for the top decile falls by roughly \$1,200 (in 1991 dollars)
Barcellos and Jacobson (2015) Medicare eligibility at age 65, U.S. adults 57–74 (HRS) At age 65, mean annual OOP medical spending falls by about 33%. The 95th percentile of the OOP distribution falls by about 53%; the share of individuals with OOP expenses exceeding income is cut by more than half, and the share with OOP spending above 50% of income also falls sharply.
Engelhardt and Gruber (2011) Medicare Part D prescription-drug benefit (2006), U.S. adults 65+ Part D yields modest or statistically weak changes in mean total OOP spending for the elderly as a whole. The policy substantially reduces high OOP drug spending: the 95th percentile and measures of “catastrophic” spending (large OOP relative to income) fall significantly, indicating improved financial protection even when mean spending moves little.
Allen et al. (2013) Medicaid lottery for low-income uninsured adults in Oregon Medicaid has modest effects on average annual OOP spending (mean levels are small in this population and imprecisely estimated). Medicaid reduces the probability of having any OOP medical spending by about 20 percentage points (a $\approx 35\%$ reduction), and lowers the probability of unpaid medical bills being sent to collection by about 6.4 percentage points ($\approx 25\%$ reduction), indicating large gains in financial protection relative to modest changes in mean spending.
Scott et al. (2021) Uninsured patients using U.S. hospital emergency departments (EDs), 2006–2017 Median ED charges for uninsured patients increase from roughly \$842 in 2006 to about \$2,033 in 2017 (in nominal terms), reflecting rising overall cost levels. The study documents a very high prevalence of catastrophic health expenditures (CHE) among uninsured ED users when full charges are compared to income, and notes prior estimates that 70–90% of uninsured ED visits resulting in hospitalization lead to CHE; the key message is the extreme risk of catastrophic burden, not a finely estimated change in mean spending.
Barcellos and Jacobson (2015), supplemental results Same HRS sample, focusing on financial strain Small or zero effects on average non-medical consumption. Large reductions in the probability of borrowing, skipping payments, or experiencing other indicators of financial strain linked to high medical bills, consistent with big risk/variance effects relative to changes in means.
Jones et al. (2018) Lifetime medical spending among U.S. retirees Mean lifetime medical spending is large, but Medicare and supplemental insurance leave expected spending relatively similar across many coverage types. The variance and upper tail of lifetime spending are very sensitive to coverage generosity; supplemental coverage mainly reduces the dispersion and extreme upper-tail risk rather than radically changing the mean.
Kluender et al. (2021) U.S. medical debt in collections, 2009–2020; ACA Medicaid expansion vs. non-expansion states In Medicaid expansion states, mean medical debt in collections per person falls by about \$400–\$450 relative to non-expansion states. The share of individuals with any medical debt in collections declines by roughly 20% in expansion states, and the upper tail of medical debt (e.g., 90th percentile) shrinks substantially, indicating large risk protection gains on top of modest average reductions.
Notes: This table summarizes a selection of studies documenting that major insurance reforms often produce relatively modest changes in mean out-of-pocket (OOP) spending but large changes in measures of financial risk and tail exposure (upper quantiles, catastrophic expenditure indicators, or debt in collections). Exact numbers are taken from the cited articles or authoritative summaries when reported; where only qualitative statements are available, we indicate that mean effects are modest or not separately reported.
Table 10: Standard errors for Q–LS coefficient under alternative methods (real OOP)
Robust OLS Bootstrap Q–LS SE
Projected real OOP (Q–LS index, 2015 k\$) $0.0072$ $0.0075$ $0.0072$
Notes: This table reports alternative standard error estimates for the Q–LS coefficient on projected real OOP spending (in thousands of 2015 dollars) from the specification in Table 7. “Robust OLS” treats $x_{\mathrm{QLS}}$ as a generated regressor and uses heteroskedasticity-robust standard errors clustered at the household level. “Bootstrap” reports the standard deviation of the coefficient across 999 bootstrap replications clustered by household. “ Q–LS SE” implements the analytical variance formula derived in Section 4, with the same clustering. All three methods yield very similar standard errors.

### v2

@@
 Our first contribution is to show that distributional relevance is a meaningful notion of instrument strength, not just a descriptive property.
-Within a nonseparable triangular model, we characterize a class of purely distributional instruments: variables $Z$ for which $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ but $F_{X\mid Z}(\cdot\mid Z)$ is non-degenerate in $L^{2}$.
+Within a non-separable triangular model, we characterize a class of purely distributional instruments: variables $Z$ for which $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ but $F_{X\mid Z}(\cdot\mid Z)$ is non-degenerate in $L^{2}$.
 Building on the control-function framework of Chernozhukov et al.
@@
 Our second, and main, contribution is to develop a practical IV estimator tailored to these designs, which we call Quantile Least Squares (Q–LS).
+[^4] Our notion of “distributional instruments” is orthogonal to the shift-share (Bartik) literature (see (Adao et al., 2019; Borusyak et al., 2025; Goldsmith-Pinkham et al., 2020) for a detailed review).
+There, the focus is on constructing a composite instrument from many shocks and exposure shares.
+Here, we take the instrument $Z$ as given and show how to optimally exploit its distributional impact on $X$ via conditional quantiles.
+In particular, Q–LS does not impose a shift–share structure on $Z$; it can be applied whether $Z$ is a simple policy indicator or a more complex shock.
 Instead of summarizing the first stage with a conditional mean, Q–LS builds instruments from the entire conditional quantile process of $X$ given $Z$.
@@
 This is the nonlinear analogue of weak instruments: the conditional distribution of $X$ given $Z$ is only weakly sensitive to $Z$, so the control function has little effective variation.
+Closest in spirit to our work is the recent “Distributional Instrumental Variable (DIV)” method of Holovchak et al.
+(2025).
+They also exploit distributional shifts in $X\mid Z$ and show how to identify and estimate the entire interventional distribution of the outcome using flexible generative models in a nonlinear IV setting.
+DIV provides conditions under which the interventional distribution is identified, and illustrates designs where 2SLS fails but distributional information suffices for identification and accurate estimation of mean and quantile treatment effects.
+Relative to this approach, we focus on average structural effects in a triangular model and on the construction of an optimal scalar instrument for linear IV estimands.
+Our Q–LS procedure can be viewed as a low-dimensional, analytically tractable counterpart to DIV: it concentrates the distributional information in a finite-dimensional quantile dictionary and delivers an optimal $L^{2}$-projection that can be used within the classical IV toolkit.
+Moreover, when the endogenous regressor is (or can be reduced to) a binary treatment, the Q–LS index provides a scalar monotone instrument, so that the resulting 2SLS coefficient admits a standard LATE interpretation as a positive-weighted average of complier effects across index increments.
 To connect with the more familiar linear IV setting, consider the linear special case
@@
 Classical weak-instrument problems arise when the first-stage coefficients $\pi$ are small, so that the concentration parameter is close to zero and the instruments explain only a tiny fraction of the variation in $X$; see Staiger and Stock (1997), Stock and Yogo (2005), and Dufour (1997) for canonical analyses.
-In this case, 2SLS and related estimators exhibit large finite-sample bias toward the OLS estimand, heavy-tailed or multimodal sampling distributions, and nonstandard asymptotic behavior ((Andrews and Mikusheva, 2014; Andrews et al., 2019) for a general overview).
+In this case, 2SLS and related estimators exhibit large finite-sample bias toward the OLS estimand, heavy-tailed or multimodal sampling distributions, and nonstandard asymptotic behavior (see Andrews and Mikusheva, 2014; Andrews et al., 2019, for a general overview).
 These issues are amplified in many-instrument settings (e.g.
-Bekker, 1994; Hansen et al., 2008), where the effective first-stage signal per instrument becomes small even if the joint $F$ is sizable.
+Bekker, 1994; Hansen et al., 2008), where the effective first-stage signal per instrument becomes small even if the joint $F$-statistic is sizable.
 The weak-identification phenomena in triangular models are conceptually similar but operate at the level of distributions rather than means.
@@
 In such cases, Q–LS leverages distributional variation in $X\mid Z$ to construct an effective instrument $h(Z)$ that satisfies a strong-IV condition even when the original mean-based instrument fails conventional first-stage diagnostics.
-Our notion of distributional relevance can be viewed as specifying a “strong-instrument” regime for triangular models at the distributional level: it requires that $F_{X\mid Z}(\cdot\mid z)$ varies nontrivially with $Z$ in an $L^{2}$ sense, even when the conditional mean $\mathbb{E}[X\mid Z]$ is flat.
+Our notion of distributional relevance thus specifies a “strong-instrument” regime at the distributional level and complements both the generative DIV approach and the classical weak-IV toolkit.
 This allows us to focus on designs where instruments may be weak for linear first stages but remain informative about the distribution of $X$, and hence can still support identification and efficient estimation once the first stage is recast in distributional terms.
@@
 then the plug-in estimator that regresses $Y$ on $(\widehat{X}^{\mathrm{QLS}},Z_{1})$ is identical (in finite samples) to the corresponding Q–LS IV estimator.
-General comparison
+General comparison and LATE interpretation
 Outside of the pure location case, Q–LS remains consistent under distributional relevance even when mean relevance fails or the conditional mean is misspecified, because it exploits shifts in the entire distribution of $X\mid Z$.
-In settings where both mean relevance and distributional relevance hold and the efficient IV $h^{*}(Z)$ lies in the closure of the quantile-generated class $\mathcal{H}$, Q–LS again attains the same efficiency bound as optimal 2SLS.
-In more general cases where $h^{*}(Z)\notin\mathcal{H}$, Q–LS is efficient within $\mathcal{H}$ but may be slightly less efficient than an oracle 2SLS that can use arbitrary functions of $Z$ as instruments.
+In designs where both mean relevance and distributional relevance hold and the efficient IV $h^{*}(Z)$ lies in the closure of the quantile-generated class $\mathcal{H}$, Q–LS again attains the same efficiency bound as optimal 2SLS.
+When $h^{*}(Z)\notin\mathcal{H}$, Q–LS is efficient within $\mathcal{H}$ but may be slightly less efficient than an oracle 2SLS that can use arbitrary functions of $Z$ as instruments.
 The main gain of Q–LS is robustness: it can remain strongly identified in settings where the classical first-stage mean shift is close to zero but the distribution of $X\mid Z$ varies substantially with $Z$.
+When the endogenous regressor is (or can be reduced to) a binary treatment $X=D\in\{0,1\}$, Q–LS also admits a familiar LATE interpretation under additional structure.
+In the positive-normalized version, the Q–LS index $H=h(Z)$ is a convex combination of conditional quantiles of $D\mid Z$, which collapses possibly high-dimensional instruments into a single scalar encouragement variable.
+Under standard IV validity and a no-defiers monotonicity condition with respect to $H$, the 2SLS coefficient using $H$ as the sole instrument can be written as a positive-weighted average of local average treatment effects for units whose treatment status changes as $H$ increases.
+Thus, in binary-treatment applications, Q–LS can be viewed not only as a strength/efficiency device but also as a way to construct a scalar monotone instrument that delivers a conventional LATE interpretation (see Appendix B for details).
+In the binary-treatment case, the scalar Q–LS index also admits a LATE representation with nonnegative weights on adjacent “index-complier” groups (Appendix B), and these weights can be estimated in practice, so the researcher can diagnose which parts of the instrument support contribute most to the Q–LS effect.
+In addition, because the Q–LS index is built as a convex combination of conditional quantiles, the estimated weight vector $\hat{w}$ reveals which parts of the $D\mid Z$ distribution (e.g.
+central vs.
+tail quantiles) contribute most to instrument strength.
 5 Implementation and regularization
@@
 A moderate level of penalization, combined with post-LASSO refitting, tends to yield stable first stages and good finite-sample size for Q–LS inference.
+5.3 Positive normalized Q–LS weights
+Our baseline Q–LS instrument is a linear functional of the conditional quantile process,
+$h_{\omega}(Z)=\int_{0}^{1}\omega(\tau)\,Q_{X\mid Z}(\tau\mid Z)\,d\tau,$
+with $\omega\in L^{2}(0,1)$.
+In finite samples, we approximate this object using a quantile grid $0<\tau_{1}<\cdots<\tau_{K}<1$ and fitted quantiles $\hat{g}_{k}(Z)\approx\widehat{Q}_{X\mid Z}(\tau_{k}\mid Z)$, so that $h(Z)\approx\hat{G}(Z)^{\prime}w$ for $\hat{G}(Z)=(\hat{g}_{1}(Z),\ldots,\hat{g}_{K}(Z))^{\prime}$.
+For interpretability and stability, we also consider a constrained version that enforces positive, normalized weights:
+$w_{\mathrm{PN}}\in\arg\min_{w\in\mathbb{R}^{K}}\ \mathbb{E}\!\left[(X-\hat{G}(Z)^{\prime}w)^{2}\right]\quad\text{s.t.}\quad w_{k}\geq 0\ \ \forall k,\qquad\mathbf{1}^{\prime}w=1.$ (22)
+Equivalently, in population form, this corresponds to restricting $\omega(\tau)\geq 0$ and $\int_{0}^{1}\omega(\tau)\,d\tau=1$.
+Interpretation.
+Under these constraints, $h_{\mathrm{PN}}(Z)=\hat{G}(Z)^{\prime}w_{\mathrm{PN}}$ is an L-statistic: it is a convex combination of conditional quantiles, i.e.
+a “quantile average” index.
+Positivity also preserves distributional ordering: if the conditional distribution of $X$ under $Z=z_{2}$ first-order stochastically dominates that under $Z=z_{1}$ (so that $Q_{X\mid Z=z_{2}}(\tau)\geq Q_{X\mid Z=z_{1}}(\tau)$ for all $\tau$), then $h_{\mathrm{PN}}(z_{2})\geq h_{\mathrm{PN}}(z_{1})$ for any admissible weight function.
+This rules out sign-cancellation across quantiles and yields an index whose movements track distributional shifts in a single direction.
+Finite-sample stability.
+Because extreme quantiles are noisier to estimate, unconstrained Q–LS may assign large positive and negative weights that amplify estimation error.
+The simplex constraint $w\geq 0$, $\mathbf{1}^{\prime}w=1$ regularizes the instrument, bounding its variability and often improving robustness when $\hat{G}$ is generated (e.g.
+via quantile regression).
+In our empirical implementation we compute $w_{\mathrm{PN}}$ by quadratic programming, and we can combine it with cross-fitting to further limit overfitting of $\hat{G}$.
 5.4 Practical recommendations
@@
 (A) strong mean relevance (benchmark Gaussian case), (B) Trigonometric and distributional instruments (variance shifts) and (C) mixed mean and distributional relevance.
-The following parts of the data generating processes (DGPs) are the same in all three designs;
+The following parts of the data-generating processes (DGPs) are the same in all three designs;
   • The structural equation
@@
 The full results, including average asymptotic standard error and simulation standard deviation, and Q–LS-L2 are available upon request.
-Table 2 presents the results for Design A and Design C for sample sizes of 500 and 1000.555Full results for Design A and Design C are available upon request.
+Table 2 presents the results for Design A and Design C for sample sizes of 500 and 1000.666Full results for Design A and Design C are available upon request.
 It is important to note that classical IV is correctly specified, as $Z_{2}$ is strongly mean-relevant in both Design A and Design C.
@@
 (2020) for related results.
-Appendix B Detailed Proofs of Lemma and Propositions
+Appendix B A LATE interpretation for Q–LS under a binary treatment
+This appendix provides a potential-outcome interpretation of Q–LS as a LATE estimator under additional structure.
+The main text treats Q–LS primarily as a strength/efficiency device for IV.
+Here we show that when the endogenous variable is binary (or can be interpreted as a binary treatment derived from a latent index), Q–LS naturally delivers a scalar monotone instrument for which the standard LATE logic applies.
+B.1 Setup: binary treatment, scalar instrument, and a Q–LS index
+Let $D\in\{0,1\}$ be the endogenous treatment and let $Y(d)$ denote potential outcomes for $d\in\{0,1\}$.
+We observe $Y=Y(D)$.
+For the LATE interpretation in this appendix, we specialize to a scalar instrument $Z$ (which can be binary, $Z\in\{0,1\}$, or more generally ordered).
+If exogenous covariates $Z_{1}$ are present, we work after partialling them out and suppress $Z_{1}$ in the notation.
+In the quantile-dictionary implementation of Q–LS, we consider a finite grid $\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}\subset(0,1)$ and define
+$G(Z):=\big(Q_{D\mid Z}(\tau_{1}\mid Z),\ldots,Q_{D\mid Z}(\tau_{K}\mid Z)\big)^{\prime},$
+where $Q_{D\mid Z}(\tau\mid Z)$ denotes the conditional $\tau$-quantile of $D$ given $Z$. Q–LS then forms scalar indices of the form
+$H_{w}(Z):=G(Z)^{\prime}w,\qquad w\in\Delta_{K}:=\big\{w\in\mathbb{R}^{K}:\ w_{k}\geq 0,\ \mathbf{1}^{\prime}w=1\big\}.$
+When $w\in\Delta_{K}$, $H_{w}(Z)$ is a convex combination of conditional quantiles—an L-statistic—which is useful both for monotonicity and for interpretability.
+Assumption A.4 (IV validity).
+(Independence and exclusion) The instrument $Z$ is independent of the potential outcomes and potential treatments, and affects $Y$ only through $D$:
+$\{Y(0),Y(1),D(z)\}_{z}\;\perp\!\!\!\perp\;Z.$
+(If covariates $Z_{1}$ are included, the independence is conditional on $Z_{1}$.)
+Assumption A.5 (Quantile relevance).
+(Distributional relevance of the scalar instrument) There exists at least one $\tau_{k}\in\mathcal{T}$ such that the conditional quantile $Q_{D\mid Z}(\tau_{k}\mid Z)$ is not almost surely constant in $Z$.
+Equivalently, not all components of $G(Z)$ are almost surely flat in $Z$.
+Remark 1.
+When $D$ is binary, the conditional quantiles $Q_{D\mid Z}(\tau\mid Z)$ are step functions of the propensity score $p(Z):=\Pr(D=1\mid Z)$.
+In this case, Assumption A.5 is equivalent to $p(Z)$ not being almost surely constant.
+We formulate relevance in terms of conditional quantiles to keep the connection with the distributional-instrument framework used in the main text.
+Quantile relevance ensures that there exist weights $w\in\Delta_{K}$ for which $\mathrm{Cov}\big(H_{w}(Z),D\big)\neq 0$, i.e. the scalar index $H_{w}$ is a relevant instrument for $D$.
+Assumption A.6 (Quantile monotonicity in the scalar instrument).
+For each $\tau\in\mathcal{T}$, the conditional quantile $z\mapsto Q_{D\mid Z}(\tau\mid z)$ is weakly increasing in $z$.
+Lemma A.1 (Convex combinations of quantiles preserve monotonicity).
+Under Assumption A.6, any index $H_{w}(Z)=G(Z)^{\prime}w$ with $w\in\Delta_{K}$ is weakly increasing in the scalar instrument $Z$.
+Proof.
+For each $\tau_{k}\in\mathcal{T}$, $z\mapsto Q_{D\mid Z}(\tau_{k}\mid z)$ is weakly increasing by Assumption A.6.
+A convex combination of weakly increasing functions is weakly increasing, so $z\mapsto H_{w}(z)=\sum_{k=1}^{K}w_{k}Q_{D\mid Z}(\tau_{k}\mid z)$ is also weakly increasing.
+∎
+In what follows, we fix a weight vector $w^{\star}\in\Delta_{K}$ (e.g. the one selected by Q–LS) that satisfies Assumption A.5, and write $H:=H_{w^{\star}}(Z)$.
+Assumption A.7 (Monotone treatment selection w.r.t. the index).
+(No defiers in the index ordering) There exists a version of the potential treatment process $\{D(h)\}_{h}$ such that
+$h_{1}\geq h_{0}\quad\Longrightarrow\quad D(h_{1})\geq D(h_{0})\qquad\text{a.s.}$
+In particular, if $H_{1}\geq H_{0}$ then $D(H_{1})\geq D(H_{0})$ almost surely.
+Assumption A.7 is the usual no-defiers (monotonicity) condition, expressed with respect to the scalar Q–LS index $H$.
+Assumption A.6 and Lemma A.1 justify interpreting $H$ as a one-dimensional “encouragement" variable inherited from the scalar instrument $Z$.
+B.2 Wald LATE using the Q–LS index
+When $H$ takes at least two values $h_{0}<h_{1}$ with positive probability, the standard Wald ratio based on those two values identifies a LATE for compliers defined relative to $H$.
+Proposition A.1 (LATE for two index values).
+Suppose Assumptions A.4, A.5, A.6, and A.7 hold.
+Let $H$ take the values $h_{0}<h_{1}$ in its support with $\Pr(H=h_{j})>0$ for $j\in\{0,1\}$ and $\mathbb{E}[D\mid H=h_{1}]\neq\mathbb{E}[D\mid H=h_{0}]$.
+Then
+$\frac{\mathbb{E}[Y\mid H=h_{1}]-\mathbb{E}[Y\mid H=h_{0}]}{\mathbb{E}[D\mid H=h_{1}]-\mathbb{E}[D\mid H=h_{0}]}=\mathbb{E}\!\left[Y(1)-Y(0)\ \big|\ D(h_{1})>D(h_{0})\right].$ (38)
+Proof.
+Under IV validity, $\mathbb{E}[Y\mid H=h]=\mathbb{E}[Y(D(h))]$ and $\mathbb{E}[D\mid H=h]=\mathbb{E}[D(h)]$.
+Write $Y(D(h))=Y(0)+D(h)\{Y(1)-Y(0)\}$ and take differences between $h_{1}$ and $h_{0}$.
+Monotonicity implies $D(h_{1})-D(h_{0})\in\{0,1\}$, which delivers (38).
+∎
+B.3 2SLS with a multi-valued scalar index: positive LATE weights
+In practice, Q–LS produces a scalar index $H$ with many support points, and empirical work typically uses 2SLS with $H$ (possibly along with controls).
+While multiple instruments can generate non-convex (possibly negative) weights across different LATEs, a key advantage of using a single scalar index is that the IV estimand can be expressed as a positive-weighted average of LATEs across index increments.
+Assume $H$ has finite support $\{h_{1}<\cdots<h_{J}\}$ for exposition, and define adjacent-index LATEs
+$\text{LATE}_{j}:=\mathbb{E}\!\left[Y(1)-Y(0)\ \big|\ D(h_{j+1})>D(h_{j})\right],\qquad j=1,\ldots,J-1.$
+Let $\Delta_{j}:=\mathbb{E}[D\mid H=h_{j+1}]-\mathbb{E}[D\mid H=h_{j}]\geq 0$ under Assumption A.7.
+Proposition A.2 (2SLS as a positive average of adjacent LATEs).
+Suppose Assumptions A.4, A.5, A.6, and A.7 hold and $H$ takes values $\{h_{1}<\cdots<h_{J}\}$ with positive probability. Then the population IV coefficient using the single instrument $H$,
+$\beta_{IV}(H):=\frac{\mathrm{Cov}(H,Y)}{\mathrm{Cov}(H,D)},$
+can be written as
+$\beta_{IV}(H)=\sum_{j=1}^{J-1}\delta_{j}\,\text{LATE}_{j},\qquad\delta_{j}:=\frac{\pi_{j}}{\sum_{\ell=1}^{J-1}\pi_{\ell}},\qquad\pi_{j}\geq 0,$
+where $\text{LATE}_{j}$ is the adjacent-index LATE between $h_{j}$ and $h_{j+1}$. In particular, the weights $\delta_{j}$ are nonnegative and sum to one.
+Remark.
+The exact form of $\pi_{j}$ depends on the distribution of $H$ (and on whether $H$ is centered), but the key point is conceptual: with a single scalar index satisfying monotonicity, the IV estimand aggregates local LATEs using positive weights tied to how much the index shifts treatment take-up across adjacent levels.
+B.4 Why Q–LS is helpful for a LATE interpretation
+Propositions A.1A.2 show that the LATE interpretation hinges on (i) validity of the underlying instrument and (ii) monotonicity with respect to a scalar index.
+Q–LS contributes by constructing such a scalar index from distributional features of the first stage.
+(i) Scalar reduction.
+Even when the original instrument vector is high-dimensional, the quantile dictionary $G(Z)$ and the Q–LS weights $w^{\star}$ collapse information into a one-dimensional index $H=H_{w^{\star}}(Z)$.
+This facilitates monotonicity assumptions and avoids the negative-weight issues that can arise when combining multiple instruments in 2SLS.
+(ii) Positive normalized weights and monotone distributional shifts.
+When $w\in\Delta_{K}$, $H_{w}(Z)$ is a convex combination of conditional quantiles.
+Under Assumption A.6, higher values of the scalar instrument $Z$ shift each conditional quantile of $D$ weakly to the right, so $H_{w}(Z)$ inherits this ordering (Lemma A.1).
+This provides a transparent route for arguing that $H$ acts as a monotone “encouragement” variable, supporting Assumption A.7.
+(iii) Strength.
+Within the class of admissible indices (e.g.
+$w\in\Delta_{K}$), Q–LS chooses weights to maximize predictive content for the first stage.
+Thus, Q–LS does not change the causal content of IV validity, but can materially improve finite-sample performance by increasing $\mathrm{Cov}(H,D)$.
+Remark 2 (Inspecting which complier groups matter most).
+The representation in Proposition A.2 is not only conceptual: in empirical work the weights $\omega_{j}$ can be estimated from the data, because they are proportional to the contribution of each adjacent index increment to $Cov(H,D)$.
+Concretely, letting $\hat{\Delta}_{j}:=\mathbb{E}_{n}[D\mid H=h_{j+1}]-\mathbb{E}_{n}[D\mid H=h_{j}]$ and $\hat{p}_{j}:=\Pr_{n}(H\in\{h_{j},h_{j+1}\})$, we can form empirical analogues $\hat{\pi}_{j}$ and $\hat{\delta}_{j}$ and report them as a diagnostic.
+This makes it possible to see which “complier bands” between $h_{j}$ and $h_{j+1}$ contribute most to the overall Q–LS estimate.
+Remark 3 (Which parts of the distribution drive the Q–LS index?).
+Because the Q–LS index takes the form $H_{w}(Z)=G(Z)^{\prime}w$ with $G(Z)=\big(Q_{X\mid Z}(\tau_{1}\mid Z),\ldots,Q_{X\mid Z}(\tau_{K}\mid Z)\big)^{\prime}$ and $w\in\Delta_{K}$, the estimated weights $\hat{w}$ provide a direct diagnostic of which quantile components are most influential in the first stage.
+Large weights on upper quantiles (e.g.
+$\tau\geq 0.8$) indicate that Q–LS is mainly leveraging tail risk, whereas mass on intermediate or lower quantiles reflects sensitivity of the effect to central or lower parts of the $X\mid Z$ distribution.
+In applications, reporting $\hat{w}$ (or aggregations such as the total weight above a given quantile) helps interpret whether the Q–LS instrument is primarily a location, dispersion, or tail-risk device.
+B.5 Scope and limitations
+The LATE interpretation in this appendix requires a binary treatment (or a binary treatment derived from a latent/continuous exposure) and monotonicity with respect to the scalar index $H$.
+When the endogenous regressor is genuinely continuous, the analogous interpretation is in terms of weighted averages of local causal responses (e.g.
+“marginal treatment effects” or “local IV” objects), rather than LATE.
+In that case, Q–LS still operates as an instrument-construction/strength device, but the exact causal interpretation is no longer a complier LATE.
+Appendix C Detailed Proofs of Lemma and Propositions
 This appendix collects the main proofs and clarifies the role of the Q–LS instrument as an $L^{2}$ projection.
 Throughout we use the same notation and assumptions as in the main text and treat the number of quantile basis functions $J$ as fixed (or growing slowly enough with $n$ that standard sieve arguments apply).
-B.1 Setup and assumptions
+C.1 Setup and assumptions
 Let $\{(Y_{i},X_{i},Z_{i})\}_{i=1}^{n}$ be i.i.d. observations, where
@@
 Assumption B.3 is the usual sieve approximation condition: as the dictionary grows (via a finer quantile grid and more basis terms), linear combinations of $G_{n}(Z_{i})$ approximate the optimal projection $h^{\star}$ in mean square.
-B.2 Instrument relevance of the optimal projection
+C.2 Instrument relevance of the optimal projection
 We first show that the $L^{2}$-projection $h^{\star}$ is a relevant instrument whenever the dictionary is distributionally relevant.
@@
 3).
-B.3 Equivalence of 2SLS and OLS with projected regressors
+C.3 Equivalence of 2SLS and OLS with projected regressors
 We next recall the standard algebra showing that 2SLS with instruments $M_{i}$ is equivalent to OLS on the regressors projected onto the instrument space.
@@
 This is the usual equivalence between 2SLS and OLS on the fitted values of the endogenous regressors; see, e.g., Wooldridge (2010) and Angrist and Imbens (1995).
-B.4 Definition and consistency of the Q–LS estimator
+C.4 Definition and consistency of the Q–LS estimator
 In the Q–LS procedure we choose a weight vector $w_{n}\in\mathbb{R}^{J_{n}}$ that makes the projected instrument $G_{n}(Z_{i})^{\prime}w_{n}$ as close as possible, in moment terms, to the optimal projection $h^{\star}(Z_{i})$.
@@
 ∎
-B.5 Asymptotic normality of Q–LS
+C.5 Asymptotic normality of Q–LS
 We now derive the asymptotic distribution of $\hat{\beta}_{n}$ under a fixed dictionary size $J_{n}=J$; extending to slowly growing $J_{n}$ requires additional notation but follows the same steps.
@@
 In practice, we estimate the asymptotic variance with the usual cluster-robust IV sandwich estimator, replacing $h^{\star}$ by $\hat{h}_{n}$ and sample expectations by empirical averages.
-B.6 Equivalence with linear 2SLS under mean-shift designs
+C.6 Equivalence with linear 2SLS under mean-shift designs
 Finally, we formalize the statement that when the first stage is exactly linear in the original policy instrument and controls, and the Q–LS dictionary contains this linear space, Q–LS coincides with ordinary 2SLS.